### v1

# Superconvergent and Divergence-Free Finite Element Methods for Stokes Equation

Abstract.
Superconvergent and divergence-free finite element methods for the Stokes equation are developed. The velocity and pressure are discretized using $H(\mathrm{div})$-conforming vector elements and discontinuous piecewise polynomials. The discrete formulation employs a weak deviatoric gradient operator built with tangential-normal continuous finite elements for traceless tensors, requiring no stabilization. Optimal and superconvergent error estimates are established. The method connects to nonconforming virtual element and pseudostress-velocity-pressure mixed formulations. Numerical experiments verify the theory.
Key words and phrases:
Stokes equation, divergence-free, superconvergence, mixed finite element method
2020 Mathematics Subject Classification:
65N12; 65N22; 65N30;
The first author was supported by NSF DMS-2309777 and DMS-2309785. The second author was supported by the National Natural Science Foundation of China Project 12171300.
1. Introduction
In this work, we develop superconvergent and divergence-free finite element methods for the Stokes equation on a bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) with external force $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{d})$:
(1) $\begin{cases}-\Delta\boldsymbol{u}-\nabla p=\boldsymbol{f}&\text{in }\Omega,\\ \operatorname{div}\boldsymbol{u}=0&\text{in }\Omega,\\ \boldsymbol{u}=0&\text{on }\partial\Omega,\end{cases}$
where $\boldsymbol{u}$ and $p$ denote the velocity and pressure.
Classical stable finite element pairs for (1) include the Taylor–Hood element [50], the MINI element [3], and the nonconforming $P_{1}$$P_{0}$ element [25]. More examples and a detailed discussion are given in [10]. However, these classical pairs are not pointwise divergence-free, which weakens mass conservation and reduces pressure robustness in error estimates. A review of the importance of divergence-free constraints and their numerical implications can be found in [41].
Divergence-free finite element pairs include the Scott–Vogelius element [48], which requires special mesh conditions; smooth finite element pairs [56, 17, 18] with supersmooth degrees of freedom; conforming pairs on split meshes [6, 55, 22, 28, 32, 38]; conforming rational pairs [33, 34]; and low-order, normal-continuous nonconforming pairs [43, 49, 54]. We also refer to [8, 52, 51] for discontinuous Galerkin methods, and to [9, 20, 53, 40] for virtual element methods.
One can impose the divergence-free condition by using an $H(\operatorname{div})$-conforming finite element space for the velocity. Specifically, we use the Raviart–Thomas (RT) element space [47, 45] or the Brezzi–Douglas–Marini (BDM) element space [12, 46]:
$\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}:=\{\boldsymbol{v}_{h}\in H_{0}(\operatorname{div},\Omega):\boldsymbol{v}_{h}|_{T}\in\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x}\;\text{ for }T\in\mathcal{T}_{h}\}$
for the velocity, with integers $\ell=k$ (${\rm RT}_{k}$) or $\ell=k-1$ (${\rm BDM}_{k}$), and the discontinuous piecewise polynomial space $\mathbb{P}_{\ell}(\mathcal{T}_{h})$ for the pressure.
Since $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\not\subset H^{1}(\Omega;\mathbb{R}^{d})$, the discretization of $\Delta\boldsymbol{u}$ is nontrivial. One approach is to use discontinuous Galerkin (DG) methods for the Laplacian operator, as in [52, 23]. The DG method in [52] introduces a penalty term to ensure stability, while the hybridizable DG (HDG) method in [23] employs $k$th-order polynomial spaces for all variables. In fact, it suffices to take $p_{h}\in\mathbb{P}_{k-1}(\mathcal{T}_{h})$ when the velocity $\boldsymbol{u}_{h}\in{\rm BDM}_{k}$.
In the vorticity–velocity–pressure formulation [26, 27], the vorticity $\omega={\rm curl\,}\boldsymbol{u}$ is introduced, and
$\Delta\boldsymbol{u}={\rm curl\,}\omega-{\rm grad\,}\operatorname{div}\boldsymbol{u}={\rm curl\,}{\rm curl\,}\boldsymbol{u}-{\rm grad\,}\operatorname{div}\boldsymbol{u}$
is interpreted as the Hodge Laplacian. However, as observed in [26, 27], for no-slip Dirichlet boundary conditions ($\boldsymbol{u}=0$ on $\partial\Omega$), there are stability issues and reduced convergence rates. This was later explained in [4]: the suboptimal convergence rate comes from the lack of a proper Hilbert complex in this setting.
The lost convergence rate can be partly recovered by mesh symmetry. In the triangular MAC (TMAC) schemes for the Stokes equation [21], mass lumping is applied to the vorticity so that it can be eliminated, and the velocity and pressure convergence rates are improved by mesh symmetry. Using the element pair $\textrm{RT}_{0}$$\textrm{P}_{0}$ for velocity and pressure (i.e., $k=\ell=0$) yields
$|I_{0,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}|_{1,h}+\|Q_{0}p-p_{h}\|\lesssim h^{\min\{1,s\}}|\ln h|^{1/2},$
where $s\geq\tfrac{1}{2}$ measures mesh symmetry and $I^{\rm\operatorname{div}}_{k,\ell}$ is the interpolation operator into $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$. With the pair $\textrm{BDM}_{1}$$\textrm{P}_{0}$ (i.e., $k=1$, $\ell=0$), the TMAC scheme achieves
$|I_{1,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}|_{1,h}+\|Q_{0}p-p_{h}\|\lesssim h^{1+\min\{1/2,s\}}.$
In [30, 31], the mass-conserving mixed formulation with stresses (MCS) finite element method was developed for the pseudostress–velocity–pressure formulation, also employing $H(\operatorname{div})$-conforming velocity spaces. The stress space in [31] is slightly nonconforming, leading to the standard $h^{k}$ convergence instead of the superconvergent $h^{k+1}$ rate. In [30], a weak symmetry constraint was introduced and the stress space was enriched, yielding $h^{k+1}$ convergence for the $\mathrm{RT}_{k}$$\mathrm{P}_{k}$ pair with $k\geq 1$.
A divergence free weak virtual element method was developed in [20], which requires stabilization and does not yield superconvergence. The main idea is to introduce a piecewise vector-valued polynomial $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}:=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$ defined on faces to enforce tangential continuity. We follow this approach and improve it to a stabilization-free scheme with superconvergence. Define the weak deviatoric gradient operator element-wise by
$(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu}),\boldsymbol{\tau})_{T}=-(\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot\boldsymbol{v},\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\boldsymbol{\mu},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}$
for any $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$. The tensor space is reduced to the traceless tensor $\mathbb{T}$, since for $\boldsymbol{\sigma}:={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{tr}{\rm grad\,}\boldsymbol{u}=\operatorname{div}\boldsymbol{u}=0$.
We propose the following mixed finite element method: find $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}$, and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ with integers $k\geq 0$ and $\ell\in\{k,k-1\}$, such that
(2a) $(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h}),\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))+(\operatorname{div}\boldsymbol{v}_{h},p_{h})$ $=(\boldsymbol{f},\boldsymbol{v}_{h}),$
(2b) $(\operatorname{div}\boldsymbol{u}_{h},q_{h})$ $=0$
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}$, and $q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$. Notably, the discrete method (2) requires no additional stabilization, despite relying on the weak differential operator $\operatorname{dev}{\rm grad\,}_{w}$.
We derive superconvergent estimates for the velocity and pressure:
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|+\|Q_{\ell}p-p_{h}\|\lesssim h^{k+1}|\boldsymbol{\sigma}|_{k+1},$
where $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u},$ and $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$. Unlike the TMAC scheme [21], such superconvergence does not require the symmetry of the triangulation. It is further utilized to construct a new superconvergent approximation of $\boldsymbol{u}$.
Our method allows $\ell=0$ an important case not covered by MCS [30, 31]. We provide an illustration of the case $(k,\ell)=(0,0)$, which can be treated as a generalization of the classical MAC scheme [35] on rectangular grids to triangulations and achieve first order $h$ convergence, no loss of convergence compared with the vorticity-velocity-pressure formulation. One can use $\textrm{BDM}_{1}$$\textrm{P}_{0}$ pair, $(k,\ell)=(1,0)$, to get a second order $h^{2}$ scheme.
We propose a hybridization of (2) by relaxing the normal continuity of the velocity, which enables the choice $(k,\ell)=(0,-1)$. The velocity is approximated by the piecewise constant vector space, and the normal continuity is enforced through a nonconforming linear element space for the pressure $p_{h}$. Even in this low-order setting, the method achieves at least first-order convergence, which is a notable feature.
The remainder of this paper is organized as follows. Section 2 discusses the weak div stability and the distributional discretization of the vector Laplacian. Section 3 develops mixed finite element methods for the Stokes equation, and Section 4 presents their equivalent discrete formulations. Finally, Section 5 reports numerical experiments that confirm and illustrate the theoretical results.
2. Distributional Discretization of the Vector Laplacian
In this section, we discretize the weak divergence operator
$\operatorname{div}:L^{2}(\Omega;\mathbb{T})\to H^{-1}(\Omega;\mathbb{R}^{d}),$
and apply it to the vector Laplacian problem. We introduce weak divergence and deviatoric gradient operators, establish their inf-sup stability, and derive a superconvergent error estimate for divergence-free velocity spaces.
2.1. Notation
Let $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) be a bounded polytope with boundary $\partial\Omega$. For a bounded domain $D$ and integer $m\geq 0$, let $H^{m}(D)$ be the Sobolev space with norm $\|\cdot\|_{m,D}$ and semi-norm $|\cdot|_{m,D}$, and let $H_{0}^{m}(D)$ be the closure of $C_{0}^{\infty}(D)$ under $\|\cdot\|_{m,D}$. Set $L^{2}(D)=H^{0}(D)$ with inner product $(\cdot,\cdot)_{D}$, and write $\|\cdot\|_{0,D}$ as $\|\cdot\|_{D}$. When $D=\Omega$, the subscript $D$ is omitted. Let $h_{D}$ denote the diameter of $D$, and $\boldsymbol{n}_{\partial D}$ the unit outward normal, abbreviated as $\boldsymbol{n}$.
We use
$H(\operatorname{div},D)$ $=\{\boldsymbol{v}\in L^{2}(D;\mathbb{R}^{d}):\operatorname{div}\boldsymbol{v}\in L^{2}(D)\},$
$H_{0}(\operatorname{div},D)$ $=\{\boldsymbol{v}\in H(\operatorname{div},D):\boldsymbol{v}\cdot\boldsymbol{n}=0\text{ on }\partial D\}.$
Let $L_{0}^{2}(D)=\{q\in L^{2}(D):\int_{D}q\,\,{\rm d}x=0\}.$ For any $B\subseteq L^{2}(D)$, denote $B/\mathbb{R}=B\cap L_{0}^{2}(D)$.
Let $\{\mathcal{T}_{h}\}_{h>0}$ be a regular family of simplicial meshes of $\Omega$, where $h=\max_{T\in\mathcal{T}_{h}}h_{T}$ and $h_{T}=\operatorname{diam}(T)$. Denote by $\mathcal{F}_{h}$ and $\mathring{\mathcal{F}}_{h}$ the sets of all faces and interior faces, respectively. For each face $F\in\mathcal{F}_{h}$, fix a unit normal $\boldsymbol{n}_{F}$. For $T\in\mathcal{T}_{h}$, let $\mathcal{F}(T)$ be the set of all (d1)(d-1)-dimensional faces of $T$.
For a face $F$ and vector $\boldsymbol{v}\in\mathbb{R}^{d}$, define its tangential projection
$\Pi_{F}\boldsymbol{v}=(\boldsymbol{I}-\boldsymbol{n}_{F}\boldsymbol{n}_{F}^{\intercal})\boldsymbol{v}.$
For two adjacent elements $T^{\pm}$ sharing $F$, define the jump
$\llbracket v\rrbracket|_{F}=v^{+}\boldsymbol{n}_{F}\cdot\boldsymbol{n}_{\partial T^{+}}+v^{-}\boldsymbol{n}_{F}\cdot\boldsymbol{n}_{\partial T^{-}},$
and on a boundary face $F\subset\partial\Omega$, set $\llbracket v\rrbracket=v|_{F}\,\boldsymbol{n}_{F}\cdot\boldsymbol{n}$.
For a domain $D$ and integer $k\geq 0$, let $\mathbb{P}_{k}(D)$ be the space of polynomials on $D$ of degree $\leq k$, $\mathbb{H}_{k}(D)=\mathbb{P}_{k}(D)\setminus\mathbb{P}_{k-1}(D)$, and $Q_{k,D}$ the $L^{2}$-projection onto $\mathbb{P}_{k}(D)$. Set $\mathbb{M}=\mathbb{R}^{d\times d}$, $\mathbb{K}$ for skew-symmetric, and $\mathbb{T}$ for traceless matrices. For $\boldsymbol{\tau}\in\mathbb{M}$,
$\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\tfrac{1}{d}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I},\qquad\operatorname{skw}\boldsymbol{\tau}=\tfrac{1}{2}(\boldsymbol{\tau}-\boldsymbol{\tau}^{\intercal}).$
Define
$H^{s}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in H^{s}(T)\}=\prod_{T\in\mathcal{T}_{h}}H^{s}(T),$
$\mathbb{P}_{k}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in\mathbb{P}_{k}(T)\}=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T),$
$\mathbb{P}_{k}(\mathcal{F}_{h})=\{v\in L^{2}(\mathcal{F}_{h}):v|_{F}\in\mathbb{P}_{k}(F)\}=\prod_{F\in\mathcal{F}_{h}}\mathbb{P}_{k}(F),$
and similarly define $\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$ for interior faces only. For a space $B(D)$ or $B(\mathcal{T}_{h})$, define
$B(D;\mathbb{X})=B(D)\otimes\mathbb{X},\quad B(\mathcal{T}_{h};\mathbb{X})=B(\mathcal{T}_{h})\otimes\mathbb{X},$
where $\mathbb{X}\in\{\mathbb{R}^{d},\mathbb{M},\mathbb{T},\mathbb{K}\}$. Let $Q_{k}$ denote the $L^{2}$-projection operator from $L^{2}(\Omega)$ onto $\mathbb{P}_{k}(\mathcal{T}_{h})$; its vector- or tensor-valued version is also denoted by $Q_{k}$. We use ${\rm grad\,}_{h}$ ($\nabla_{h}$ for scalar function) and $\operatorname{div}_{h}$ to represent the element-wise gradient and divergence operators, respectively.
2.2. Weak div stability for the scalar Laplacian
We first consider the surjective map
$\operatorname{div}:L^{2}(\Omega;\mathbb{R}^{d})\to H^{-1}(\Omega),$
and discuss its distributional finite element discretization. This operator is part of the distributional de Rham complex introduced in [11].
The $L^{2}$ space is discretized by the discontinuous polynomial space
$\Sigma_{k}^{-1}(\mathcal{T}_{h};\mathbb{R}^{d})=\mathbb{P}_{k}(\mathcal{T}_{h};\mathbb{R}^{d}),$
which we abbreviate as $\Sigma_{k}^{-1}$. A natural norm on $\Sigma_{k}^{-1}$ is the standard $L^{2}$ norm $\|\cdot\|$.
Next, we introduce the product space of discontinuous polynomial spaces on $\mathcal{T}_{h}\times\mathcal{F}_{h}$:
$M_{k-1,k}^{-1}:=\mathbb{P}_{k}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathcal{F}_{h}).$
An element of $M_{k-1,k}^{-1}$ is denoted by $u=(u_{0},u_{b})$. A natural product-type inner product is defined by
$((u_{0},u_{b}),(v_{0},v_{b}))_{0,h}=(u_{0},v_{0})+\sum_{F\in\mathcal{F}_{h}}h_{F}(u_{b},v_{b})_{F},$
where the scaling factor $h_{F}$ ensures that the two terms are of comparable magnitude. The corresponding norm on $M_{k-1,k}^{-1}$ is
$\|u\|_{0,h}^{2}=\|u_{0}\|^{2}+\sum_{F\in\mathcal{F}_{h}}\|h_{F}^{1/2}u_{b}\|_{F}^{2}.$
The subspace
$\mathring{M}_{k-1,k}^{-1}=\mathbb{P}_{k}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$
can be viewed as a subspace of ${M}_{k-1,k}^{-1}$ by setting the values on boundary faces to zero.
For $\sigma\in\Sigma_{k}^{-1}$, define the operator $\operatorname{div}_{w}:\Sigma_{k}^{-1}\to\mathring{M}_{k-1,k}^{-1}$ by
$\operatorname{div}_{w}\sigma=\big(\operatorname{div}_{T}\sigma,-h_{F}^{-1}[\sigma\cdot\boldsymbol{n}]|_{F}\big)_{T\in\mathcal{T}_{h},~F\in\mathring{\mathcal{F}}_{h}}\in\mathring{M}_{k-1,k}^{-1}.$
When $\sigma\in\Sigma_{k}^{-1}\cap H(\operatorname{div},\Omega)$, we have $\operatorname{div}_{w}\sigma=(\operatorname{div}\sigma,0)$ since $\sigma\cdot\boldsymbol{n}]|_{F}=0$ for all interior faces $F$; in this case we simply write $\operatorname{div}_{w}\sigma=\operatorname{div}\sigma$.
A graph norm associated with $\operatorname{div}_{w}$ is defined as
$\|\sigma\|_{\operatorname{div}_{w}}^{2}:=\|\sigma\|^{2}+\|\operatorname{div}_{w}\sigma\|_{0,h}^{2}=\|\sigma\|^{2}+\|\operatorname{div}_{h}\sigma\|^{2}+\|h^{-1/2}[\sigma\cdot\boldsymbol{n}]\|_{\mathring{\mathcal{F}}_{h}}^{2},$
where the scaling is chosen so that the last two terms are of comparable order. For $\sigma\in\Sigma_{k}^{-1}\cap H(\operatorname{div},\Omega)$, we recover $\|\sigma\|_{\operatorname{div}_{w}}=\|\sigma\|_{\operatorname{div}}$.
The following inf-sup condition can be established in a manner similar to Theorem 3.2, and hence the proof is omitted.
Lemma 2.1.
We have the weak div stability: there exists a constant $\alpha>0$, independent of $h$, such that
(3) $\inf_{v\in\mathring{M}^{-1}_{k-1,k}}\sup_{\tau\in\Sigma_{k}^{-1}}\frac{(\operatorname{div}_{w}\tau,v)_{0,h}}{\|\tau\|_{\operatorname{div}_{w}}\|v\|_{0,h}}=\alpha>0.$
Consider the mixed formulation of the Poisson equation $\Delta u=f$ with Dirichlet boundary condition $u|_{\partial\Omega}=0$. Find $\sigma\in\Sigma_{k}^{-1}$ and $u\in\mathring{M}_{k-1,k}^{-1}$ such that
(4a) $(\sigma,\tau)$ $=(\nabla_{w}u,\tau)\quad\forall\,\tau\in\Sigma_{k}^{-1},$
(4b) $(\operatorname{div}_{w}\sigma,v)$ $=-(f,v_{0})\quad\;\forall\,v\in\mathring{M}_{k-1,k}^{-1}.$
This formulation is precisely the hybridized mixed finite element method using the ${\rm BDM}_{k}$$\mathbb{P}_{k-1}$ pair [2]. The problem (4) is well-posed by the inf-sup condition (3).
2.3. Finite element spaces
In the Stokes problem, we use the weak divergence operator to discretize the vector Laplacian. Let $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ and $\Delta\boldsymbol{u}=-\operatorname{div}\boldsymbol{\sigma}$. To distinguish the vector formulation from the scalar case, we use boldface symbols such as $\boldsymbol{\sigma}$ and $\boldsymbol{u}$.
By the tensor-product construction, we obtain the weak divergence operator
$\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{M})\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}),$
where
$\Sigma_{k}^{-1}(\mathbb{X}):=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T;\mathbb{X})\quad\text{for }\mathbb{X}=\mathbb{M}\text{ or }\mathbb{T};\quad\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}):=\mathbb{R}^{d}\otimes\mathring{M}_{k-1,k}^{-1}.$
The tensor space is restricted to the traceless subspace $\mathbb{T}$, since for $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{div}\boldsymbol{u}=0$. We then impose suitable continuity conditions on the range space and identify the subspaces of $\Sigma_{k}^{-1}(\mathbb{T})$ that lead to weakly divergence-stable pairs.
Spaces for the velocity
We employ $H(\operatorname{div})$-conforming finite elements to discretize the velocity field $\boldsymbol{u}$. For a simplex $T$ and integers $k\geq 0$ and $\ell\geq-1$, the local space of shape functions is defined by
$\mathbb{V}_{k,\ell}^{\operatorname{div}}(T):=\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x},\quad\ell=k\text{ or }k-1.$
For $\ell\geq 0$, the degrees of freedom (DoFs) are given by (cf. [16, Section 3])
(5a) $(\boldsymbol{v}\cdot\boldsymbol{n},q)_{F},$ $\quad q\in\mathbb{P}_{k}(F),\quad F\in\mathcal{F}(T),$
(5b) $(\boldsymbol{v},\boldsymbol{q})_{T},$ $\quad\boldsymbol{q}\in{\rm grad\,}\mathbb{P}_{\ell}(T)\oplus\{\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d}):\boldsymbol{q}\cdot\boldsymbol{x}=0\}.$
When $\ell=k$, the space corresponds to the Raviart–Thomas (RT) element [47, 45], where the interior moments (5b) span $\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$. For $\ell=k-1$, the Brezzi–Douglas–Marini (BDM) element [12, 46], the interior moments (5b) forms a strict subspace.
The corresponding global finite element spaces are defined as
$\mathbb{V}_{k,\ell}^{\operatorname{div}}(\mathcal{T}_{h})$ $:=\{\boldsymbol{v}_{h}\in H(\operatorname{div},\Omega):\boldsymbol{v}_{h}|_{T}\in\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)\text{ for all }T\in\mathcal{T}_{h}\},$
$\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}(\mathcal{T}_{h})$ $:=\mathbb{V}_{k,\ell}^{\operatorname{div}}(\mathcal{T}_{h})\cap H_{0}(\operatorname{div},\Omega).$
For simplicity, we omit $\mathcal{T}_{h}$ in the notation afterwards.
To study the weak div stability, we embed $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ into $\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$.
Lemma 2.2.
The mapping $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ defined by
$\boldsymbol{u}\mapsto\big(Q_{k-1,T}\boldsymbol{u},\,(\boldsymbol{0},\,\boldsymbol{n}_{F}\cdot\boldsymbol{u})_{F}\big)_{T\in\mathcal{T}_{h},\,F\in\mathring{\mathcal{F}}_{h}}$
is injective.
Proof.
If $(\boldsymbol{n}_{F}\cdot\boldsymbol{u})|_{F}=0$ and $Q_{k-1,T}\boldsymbol{u}=0$, then all degrees of freedom in (5) vanish, which implies $\boldsymbol{u}=0$. Hence, $E$ is injective. ∎
The embedding $E$ can be extended to $H^{1}(\Omega;\mathbb{R}^{d})$, but the injectivity no longer holds.
Spaces for the pseudo-stress
For an integer $k\geq 0$, the local space of shape functions is $\mathbb{P}_{k}(T;\mathbb{T})$, with degrees of freedom (DoFs) defined by
(6a) $\int_{F}\boldsymbol{t}_{i}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}\,q\,\mathrm{d}S,$ $\quad q\in\mathbb{P}_{k}(F),\ i=1,2,\ldots,d-1,\ F\in\mathcal{F}(T),$
(6b) $\int_{T}\boldsymbol{\tau}:\boldsymbol{q}\,\mathrm{d}x,$ $\quad\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{T}).$
The unisolvence of these DoFs is established in [19, 30]. The corresponding global finite element space is
[…]
Functions in $\Sigma_{k}^{\rm tn}$ are tangential-normal continuous, i.e., $\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}]_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$.
The weak divergence operator $\operatorname{div}_{w}$ defines a bilinear form on $\Sigma_{k}^{\rm tn}\times\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$:
$(\operatorname{div}_{w}\boldsymbol{\sigma},\boldsymbol{v})_{0,h}=(\operatorname{div}_{w}\boldsymbol{\sigma},E\boldsymbol{v})_{0,h}$ $=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}],\boldsymbol{n}\cdot\boldsymbol{v})_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}\boldsymbol{v})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}])_{F}.$
Remark 2.3.
The bilinear form $(\operatorname{div}_{w}\cdot,\cdot)_{0,h}$ defines a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{\rm tn}\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})^{\prime}$. With the inner product $(\cdot,\cdot)_{0,h}$, we identify $\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ with its dual. Through the embedding $E$, we further identify $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cong E(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})$ with its dual, and hence can write the mapping as
$\operatorname{div}_{w}:\Sigma_{k}^{\rm tn}\to\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}.$
2.4. Weak div stability
For each $T\in\mathcal{T}_{h}$, let $I_{T}^{\rm tn}:H^{1}(T;\mathbb{T})\to\mathbb{P}_{k}(T;\mathbb{T})$ denote the local interpolation operator defined by DoFs (6). The corresponding global interpolation operator $I_{k}^{\rm tn}:H^{1}(\mathcal{T}_{h};\mathbb{T})\to\Sigma_{k}^{-1}(\mathbb{T})$ is given by
$(I_{k}^{\rm tn}\boldsymbol{\tau})|_{T}:=I_{T}^{\rm tn}(\boldsymbol{\tau}|_{T}),\quad\forall\,T\in\mathcal{T}_{h},\ \boldsymbol{\tau}\in H^{1}(\mathcal{T}_{h};\mathbb{T}).$
Then $I_{k}^{\rm tn}\boldsymbol{\tau}\in\Sigma_{k}^{\rm tn}$ for $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$. For any $T\in\mathcal{T}_{h}$ and $\boldsymbol{\tau}\in H^{m}(T;\mathbb{T})$ with $1\leq m\leq k+1$, the standard interpolation estimate holds:
(7) $\|\boldsymbol{\tau}-I_{T}^{\rm tn}\boldsymbol{\tau}\|_{T}+h_{T}|\boldsymbol{\tau}-I_{T}^{\rm tn}\boldsymbol{\tau}|_{1,T}\lesssim h_{T}^{m}|\boldsymbol{\tau}|_{m,T}.$
We now prove the following important commutative property.
Lemma 2.4.
For $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$, we have
(8) $(\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}_{h})=(\operatorname{div}_{w}I_{k}^{\rm tn}\boldsymbol{\tau},\boldsymbol{v}_{h})_{0,h},\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.$
Proof.
Notice that $\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h}\in\Sigma_{k-1}^{-1}(\mathbb{T})$ and $\Pi_{F}\boldsymbol{v}_{h}]\in\mathbb{P}_{k}(F;\mathbb{R}^{d-1})$. So
$(\operatorname{div}_{w}I_{k}^{\rm tn}\boldsymbol{\tau},\boldsymbol{v}_{h})_{0,h}$ $=-\sum_{T\in\mathcal{T}_{h}}(I_{k}^{\rm tn}\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}I_{k}^{\rm tn}\boldsymbol{\tau}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}_{h}])_{F}$
$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\tau}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}_{h}])_{F}$
$=(\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}_{h}).$
This leads to the following inf-sup condition.
Theorem 2.5.
There exists a constant $\alpha>0$, independent of $h$, such that
(9) $\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|\boldsymbol{v}_{h}\|}=\alpha.$
Proof.
For any $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\subset L^{2}(\Omega;\mathbb{R}^{d})$, there exists $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$ such that $\operatorname{div}\boldsymbol{\tau}=\boldsymbol{v}_{h}$. Let $\boldsymbol{\tau}_{h}=I_{k}^{\rm tn}\boldsymbol{\tau}$; then the commutative property (8) implies the desired result (9). ∎
If the velocity space is enriched to ${\rm RT}_{k}$, i.e., $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$, then $\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})$. Since DoF (6b) to define $I_{k}^{\rm tn}$ is for moments of degree k1k-1, the commutative property (8) no longer holds. Consequently, the inf-sup condition may fail for the pair $\Sigma_{k}^{\rm tn}$-$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
Define the weak deviatoric gradient $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\Sigma_{k}^{\rm tn}$ as the adjoint of $\operatorname{div}_{w}$. For $\boldsymbol{v}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v}\in\Sigma_{k}^{\rm tn}$ is defined by
$(\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v},\boldsymbol{\tau})=-(\boldsymbol{v},\operatorname{div}_{w}\boldsymbol{\tau})_{0,h},\quad\forall\,\boldsymbol{\tau}\in\Sigma_{k}^{\rm tn}.$
Since $\operatorname{div}_{w}\Sigma_{k}^{\rm tn}=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, the operator $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\to\Sigma_{k}^{\rm tn}$ is injective, and thus $\|\operatorname{dev}{\rm grad\,}_{w}(\cdot)\|$ defines a norm on $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$.
The inf-sup condition (9) can then be written equivalently as
(10) $\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|\,\|\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v}_{h}\|}=1>0.$
Consider the mixed finite element method for the vector Laplacian
$\Delta\boldsymbol{u}=\boldsymbol{f},\quad\text{ with }\;\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d}).$
Find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ such that
(11) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}$ $=0,\qquad\qquad\;\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},$
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.$
The formulation (11) is well posed by the inf-sup conditions (9) or (10). However, the errors $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}I^{\operatorname{div}}_{k,k-1}\boldsymbol{u}\|$ are coupled. As a result, the dominant contribution is the $H^{1}$-type error of $\boldsymbol{u}$ of order $h^{k}$, rather than the $L^{2}$ error of $\boldsymbol{\sigma}$ of order $h^{k+1}$. This suboptimality comes from the loss of exact divergence-free conformity:
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0\ \ \forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\ \ \nRightarrow\ \ (\operatorname{div}_{w}\boldsymbol{\sigma}_{h},E\boldsymbol{v})_{0,h}=0\ \ \forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}).$
2.5. Commuting property for the weak dev grad operator
For each $T\in\mathcal{T}_{h}$, let $I_{(k,\ell),T}^{\rm div}:H^{1}(T;\mathbb{R}^{d})\to\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ denote the local interpolation operator defined by the DoFs in (5). The corresponding global interpolation operator $I_{k,\ell}^{\operatorname{div}}:H^{1}(\mathcal{T}_{h};\mathbb{R}^{d})\to L^{2}(\Omega;\mathbb{R}^{d})$ is given by
$(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v})|_{T}:=I_{(k,\ell),T}^{\rm div}(\boldsymbol{v}|_{T}),\quad\forall\,T\in\mathcal{T}_{h},\ \boldsymbol{v}\in H^{1}(\mathcal{T}_{h};\mathbb{R}^{d}).$
Then $I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\in\mathbb{V}_{k,\ell}^{\operatorname{div}}$ for any $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})$. For $T\in\mathcal{T}_{h}$ and $\boldsymbol{v}\in H^{m}(T;\mathbb{R}^{d})$ with $1\leq m\leq k+1$, the interpolation satisfies
(12) $\|\boldsymbol{v}-I_{(k,\ell),T}^{\rm div}\boldsymbol{v}\|_{T}+h_{T}|\boldsymbol{v}-I_{(k,\ell),T}^{\rm div}\boldsymbol{v}|_{1,T}\lesssim h_{T}^{m}|\boldsymbol{v}|_{m,T}.$
We are going to prove the following commutative property:
(13) $Q_{k}^{\rm tn}\operatorname{dev}{\rm grad\,}=\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}},$
where $Q_{k}^{\rm tn}:L^{2}(\Omega;\mathbb{T})\to\Sigma_{k}^{\rm tn}$ is the $L^{2}$-projection.
Lemma 2.6.
For any $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$,
(14) $(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\tau}_{h})=(\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}}\boldsymbol{u},\boldsymbol{\tau}_{h}),\quad\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}.$
Proof.
As $\operatorname{div}(\boldsymbol{\tau}_{h}|_{T})\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$ and $\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]|_{F}\in\mathbb{P}_{k}(F)$, we have
$(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\tau}_{h})$ $=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(\boldsymbol{u}\cdot\boldsymbol{n},[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}])_{F},$
$=-\sum_{T\in\mathcal{T}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}\cdot\boldsymbol{n},[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}])_{F}$
$=(\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}}\boldsymbol{u},\boldsymbol{\tau}_{h}).$
The $\mathrm{RT}_{k}$ space is required to ensure the interior moments (5b) cover $\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$ for $\operatorname{div}\boldsymbol{\tau}_{h}$, while for the $\mathrm{BDM}_{k}$ space $\mathbb{V}_{k,k-1}^{\operatorname{div}}$, the interior moments are not sufficient. Hence, the commutative property (14) does not hold for $I_{k,k-1}^{\operatorname{div}}$.
The weak operator $\operatorname{div}_{w}$ can be regarded as a discretization of $\operatorname{div}$ in the distributional sense. We can preserve the exact divergence condition in the following sense.
Lemma 2.7.
Let $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$. If
$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0,\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}},$
then $\operatorname{div}\boldsymbol{\sigma}_{h}=0$.
Proof.
It suffices to prove $\operatorname{div}\boldsymbol{\sigma}_{h}=0$ in the distributional sense:
$\langle\operatorname{div}\boldsymbol{\sigma}_{h},\boldsymbol{v}\rangle=0,\quad\forall\,\boldsymbol{v}\in C_{0}^{\infty}(\Omega;\mathbb{R}^{d}).$
Let $\boldsymbol{v}_{I}=I_{k,k}^{\operatorname{div}}\boldsymbol{v}$. Then by Lemma 2.6,
$\langle\operatorname{div}\boldsymbol{\sigma}_{h},\boldsymbol{v}\rangle=-(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}\boldsymbol{v})=-(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v}_{I})=(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{I})_{0,h}=0.$
2.6. Restriction to divergence-free subspaces
The conditions (8) and (14) are incompatible. Weak div stability holds only on the smaller space $\mathbb{V}_{k,k-1}^{\operatorname{div}}$, while the exact divergence-freeness requires the larger space $\mathbb{V}_{k,k}^{\operatorname{div}}$. We resolve this conflict by considering the divergence-free subspaces.
We have the following commuting property (cf. [5]):
(15) $\operatorname{div}\!\big(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\big)=Q_{\ell}\big(\operatorname{div}\boldsymbol{v}\big),\qquad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}),$
which implies $\operatorname{div}\!\big(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\big)=0$ for all $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$. The additional interior moments in ${\rm RT}_{k}$ compared with ${\rm BDM}_{k}$ serve only to enrich the range of the divergence operator. Hence, for any $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ and $k\geq 1$,
$I_{k,k}^{\operatorname{div}}\boldsymbol{v}=I_{k,k-1}^{\operatorname{div}}\boldsymbol{v}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
Consider the mixed finite element method for the vector Laplacian
(16) $\Delta\boldsymbol{u}=\boldsymbol{f},\quad\text{with }\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div}).$
Find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$ such that
(17a) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}$ $=0,\qquad\qquad\;\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},$
(17b) $(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
Theorem 2.8.
Let $k\geq 0$. The mixed finite element method (17) is well posed. Let $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u}$, where $\boldsymbol{u}$ solves (16). Then
(18) $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\|=\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\leq\|\boldsymbol{\sigma}-I_{k}^{\rm tn}\boldsymbol{\sigma}\|,$
and
(19) $\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|=\|Q_{k}^{\rm tn}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h})\|\leq\|\boldsymbol{\sigma}-I_{k}^{\rm tn}\boldsymbol{\sigma}\|.$
Moreover, if $\boldsymbol{u}\in H^{k+2}(\Omega;\mathbb{R}^{d})$, then
(20) $\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|+\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\|\lesssim h^{k+1}\|\boldsymbol{u}\|_{k+2}.$
Proof.
The well-posedness follows from the inf-sup condition (9) and the fact
$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}\cap\ker(\operatorname{div})=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
Denote by $\boldsymbol{\sigma}_{I}=I_{k}^{\rm tn}\boldsymbol{\sigma}$ and $\boldsymbol{u}_{I}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$. By (8), it holds that $(\operatorname{div}_{w}\boldsymbol{\sigma}_{I},\boldsymbol{v}_{h})_{0,h}=(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v}_{h})=-(\boldsymbol{f},\boldsymbol{v}_{h})=(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$, so we conclude
$(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{v}_{h})_{0,h}=0,\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
We verify the following orthogonality by using (14)
$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})$ $=(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})+(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{h})_{0,h}$
$=(\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{I},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})=-(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{I})_{0,h}=0.$
The estimate (18) follows from this orthogonality:
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|^{2}=(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\sigma}-\boldsymbol{\sigma}_{I})\leq\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{I}\|.$
By the commutative property (13), $\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{I}=Q_{k}^{\rm tn}\operatorname{dev}{\rm grad\,}\boldsymbol{u}=Q_{k}^{\rm tn}\boldsymbol{\sigma}$. Then (19) follows. ∎
The velocity space $\boldsymbol{v}_{h}|_{T}\in\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ and $\mathbb{P}_{k}(T;\mathbb{R}^{d})\subseteq\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, and the optimal order of convergence for $H^{1}$ type error is $h^{k}$. So (20) is a superconvergent error estimate for $\boldsymbol{u}$.
3. Mixed Finite Element Methods for Stokes Equation
In this section, we introduce the pressure as a Lagrange multiplier to enforce the divergence-free condition, and another multiplier to relax the tangential-normal continuity of the stress space. The resulting mixed finite element methods for the Stokes equation are pointwise divergence-free and superconvergent.
3.1. Spaces of Lagrange multipliers
The subspace $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$ is not constructed explicitly. Instead, we enforce the constraint $\operatorname{div}\boldsymbol{u}=0$ by introducing a Lagrange multiplier, which has the physical interpretation of pressure.
In the discretization, we use the piecewise polynomial space $\mathbb{P}_{\ell}(\mathcal{T}_{h})$ to approximate the pressure $p$. By the commutative property (15), it follows that
$\operatorname{div}\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}=\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}.$
We introduce the Lagrange multiplier spaces
$\Lambda_{k}=\mathbb{P}_{k}(\mathcal{F}_{h};\mathbb{R}^{d-1})\quad\textrm{and}\quad\mathring{\Lambda}_{k}=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$
to relax the tangential-normal continuity condition in the space $\Sigma_{k}^{\rm tn}$. The embedding $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ is extended as
$(\boldsymbol{u},\boldsymbol{\lambda})\mapsto\big(Q_{k-1,T}\boldsymbol{u},\,(\boldsymbol{\lambda},\,\boldsymbol{n}_{F}\cdot\boldsymbol{u})_{F}\big)_{T\in\mathcal{T}_{h},\,F\in\mathring{\mathcal{F}}_{h}},$
which remains injective.
The stress space is relaxed to $\Sigma_{k}^{-1}(\mathbb{T})=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T;\mathbb{T})$. The weak divergence operator $\operatorname{div}_{w}$ then induces the bilinear form on $\Sigma_{k}^{-1}(\mathbb{T})\times(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$:
(21) $(\operatorname{div}_{w}\boldsymbol{\sigma},(\boldsymbol{v},\boldsymbol{\mu}))_{0,h}=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}\!\Big(([\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}],\,\boldsymbol{n}\!\cdot\!\boldsymbol{v})_{F}+([\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}],\,\boldsymbol{\mu})_{F}\Big),$
which induces a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{T})\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})^{\prime}\cong(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$ by the $(\cdot,\cdot)_{0,h}$ inner product through the embedding $E$.
Its adjoint, still denoted by $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\to\Sigma_{k}^{-1}(\mathbb{T})$, is defined by
$(\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu})):=-(\operatorname{div}_{w}\boldsymbol{\sigma},(\boldsymbol{v},\boldsymbol{\mu}))_{0,h},$
for $\boldsymbol{\sigma}\in\Sigma_{k}^{-1}(\mathbb{T})$, $\boldsymbol{v}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, and $\boldsymbol{\mu}\in\mathring{\Lambda}_{k}$.
The operator $\operatorname{dev}{\rm grad\,}_{w}$ can be extended to
$\operatorname{dev}{\rm grad\,}_{w}:H^{1}(\mathcal{T}_{h};\mathbb{R}^{d})\times L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})\to\Sigma_{k}^{-1}(\mathbb{T})$
as follows: for $(\boldsymbol{v},\boldsymbol{\mu})\in H^{1}(\mathcal{T}_{h};\mathbb{R}^{d})\times L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})$, let $\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu})\in\Sigma_{k}^{-1}(\mathbb{T})$ be defined elementwise by
$(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu}),\boldsymbol{\tau})_{T}=-(\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot\boldsymbol{v},\,\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\boldsymbol{\mu},\,\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T},$
for all $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$.
Lemma 3.1.
For $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})$ and $\boldsymbol{\mu}\in L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})$, it holds that
(22) $\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{v},\,Q_{k,\mathcal{F}_{h}}\boldsymbol{\mu})=\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\,\boldsymbol{\mu}),$
where $Q_{k,\mathcal{F}_{h}}$ is the $L^{2}$-orthogonal projector from $L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})$ onto $\Lambda_{k}$.
Proof.
By the definition of the operator $\operatorname{dev}{\rm grad\,}_{w}$, we have for any $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$ that
$\quad\;(\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{v},Q_{k,\mathcal{F}_{h}}\boldsymbol{\mu}),\boldsymbol{\tau})_{T}$
$=-(I_{k,k}^{\operatorname{div}}\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot(I_{k,k}^{\operatorname{div}}\boldsymbol{v}),\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\boldsymbol{\mu},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}.$
Then (22) follows from the definition of the interpolation operator $I_{k,k}^{\operatorname{div}}$. ∎
3.2. Weak div stability
We first establish the following inf-sup condition.
Theorem 3.2.
There exists $\alpha>0$, independent of $h$, such that
(23) $\inf_{(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\times\mathring{\Lambda}_{k}}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})\|_{0,h}}=\alpha.$
Proof.
Given $\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}$, construct $\boldsymbol{\tau}_{1}\in\Sigma_{k}^{-1}(\mathbb{T})$ elementwise by
$\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}=-\tfrac{1}{2}h_{F}\,\boldsymbol{\mu}_{h}\ \ \text{on }F\in\mathcal{F}(T),\qquad(\boldsymbol{\tau}_{1},\boldsymbol{q})_{T}=0\ \ \forall\,\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{T}),$
for each $T\in\mathcal{T}_{h}$, where $\boldsymbol{n}$ is the unit outward normal on $\partial T$. Then $\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}]=h_{F}\boldsymbol{\mu}_{h}$. By inverse and scaling estimates,
$\|\boldsymbol{\tau}_{1}\|_{\operatorname{div}_{w}}\lesssim\|h_{F}^{1/2}\boldsymbol{\mu}_{h}\|_{\mathcal{F}_{h}}=\|(0,\boldsymbol{\mu}_{h})\|_{0,h}.$
By the definition of $\operatorname{div}_{w}$,
$(\operatorname{div}_{w}\boldsymbol{\tau}_{1},(0,\boldsymbol{\mu}_{h}))_{0,h}=-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}],\boldsymbol{\mu}_{h})_{F}=\|(0,\boldsymbol{\mu}_{h})\|_{0,h}^{2},$
and hence
$\|(0,\boldsymbol{\mu}_{h})\|_{0,h}\lesssim\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},(0,\boldsymbol{\mu}_{h}))_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}}\qquad\forall\,\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}.$
On the other hand, by the weak divergence inf-sup (cf. (9)),
$\|(\boldsymbol{v}_{h},0)\|_{0,h}=\|\boldsymbol{v}_{h}\|\lesssim\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}}\qquad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}},\ \boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}.$
Combining the two bounds yields (23). ∎
We use $(\cdot,\cdot)_{0,h}$ as a Riesz map to identify $E(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$ with its dual. The bilinear form (21) thus induces a mapping
$\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{T})\to\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}.$
The inf-sup condition (23) implies $\operatorname{div}_{w}$ is surjective. Consequently the adjoint $\operatorname{dev}{\rm grad\,}_{w}$ is injective and
$\|\operatorname{dev}{\rm grad\,}_{w}(\cdot,\cdot)\|\text{ defines a norm on }\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\times\mathring{\Lambda}_{k}.$
Notice that the velocity space cannot be enlarged to $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$. But later on we will use the coercivity on the divergence free subspaces.
3.3. Mixed finite element methods
With the operator $\operatorname{dev}{\rm grad\,}_{w}$, we formulate the following mixed finite element method for the Stokes equation (1): Find $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}$, and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$, for integers $k\geq 0$ and $\ell\in\{k,\,k-1\}$, such that
(24a) $(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h}),\,\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))+(\operatorname{div}\boldsymbol{v}_{h},\,p_{h})$ $=(\boldsymbol{f},\,\boldsymbol{v}_{h}),$
(24b) $(\operatorname{div}\boldsymbol{u}_{h},\,q_{h})$ $=0,$
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}$, and $q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$.
Theorem 3.3.
The mixed finite element method (24) is well-posed, with a unique solution $(\boldsymbol{u}_{h},\,\boldsymbol{\lambda}_{h},\,p_{h})\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\times\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$. Let $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$; then $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$ satisfy the mixed formulation (17).
Proof.
We first show that the mixed method (24) admits only the zero solution when $\boldsymbol{f}=0$. From (24b), it follows that $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$. Taking $\boldsymbol{v}_{h}=\boldsymbol{u}_{h}$ and $\boldsymbol{\mu}_{h}=\boldsymbol{\lambda}_{h}$ in (24a), we obtain $\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})=0$. Combining this with (23) gives $\boldsymbol{u}_{h}=0$ and $\boldsymbol{\lambda}_{h}=0$. Furthermore, by $\operatorname{div}\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}=\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$, equation (24a) implies $p_{h}=0$. Therefore, the mixed finite element method (24) is well-posed.
Since $\operatorname{div}\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}=\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$, equation (24b) implies that $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$. Restricting $\boldsymbol{v}_{h}$ to this subspace, equation (24a) becomes
(25) $(\boldsymbol{\sigma}_{h},\,\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))=(\boldsymbol{f},\,\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div}),\ \boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}.$
Taking $\boldsymbol{v}_{h}=0$ in (25) yields $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$. Then (25) reduces to equation (17b), while equation (17a) follows directly from the definition of $\boldsymbol{\sigma}_{h}$. ∎
We circumvent the inf-sup condition, as the error analysis will be derived through the equivalence to (17) without relying on the inf-sup constant.
3.4. Error analysis
Hereafter, let $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$ denote the solution of the Stokes equation (1), and set $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}=\operatorname{dev}{\rm grad\,}\boldsymbol{u}$. Then the first equation in (1) becomes
(26) $\operatorname{div}\boldsymbol{\sigma}-\nabla p=\boldsymbol{f}\quad\text{ in }\Omega.$
Moreover, let $\boldsymbol{\lambda}\in L^{2}(\mathcal{F}_{h};\mathbb{R}^{d-1})$ be defined facewise by
$\boldsymbol{\lambda}|_{F}=\Pi_{F}\boldsymbol{u},\quad\forall\,F\in\mathcal{F}_{h}.$
This notation will be used throughout the subsequent analysis.
Theorem 3.4.
Let $(\boldsymbol{u},p)$ denote the solution of the Stokes equation (1), and let $(\boldsymbol{u}_{h}$, $\boldsymbol{\lambda}_{h}$, $p_{h})$ be the solution of the mixed finite element method (24). Assume $\boldsymbol{u}\in H^{k+2}(\Omega;\mathbb{R}^{d})$. Then
(27) $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}+\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}\boldsymbol{\lambda}-\boldsymbol{\lambda}_{h})\|$
$+\|Q_{\ell}p-p_{h}\|$ $\lesssim h^{k+1}|\boldsymbol{u}|_{k+2},$
where $\|\boldsymbol{\tau}\|_{0,h}^{2}:=\|\boldsymbol{\tau}\|^{2}+\sum_{F\in\mathcal{F}_{h}}h_{F}\|\Pi_{F}\boldsymbol{\tau}\boldsymbol{n}\|_{F}^{2}$.
Proof.
By the equivalence of (24) and (17), the estimate of $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ follows from (20) in Theorem 2.8.
Then by Lemma 3.1,
$\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u},Q_{k,\mathcal{F}_{h}}\boldsymbol{\lambda})$ $=Q_{k}\operatorname{dev}{\rm grad\,}\boldsymbol{u}=Q_{k}\boldsymbol{\sigma},$
$\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$ $=\boldsymbol{\sigma}_{h}.$
So
$\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}\boldsymbol{\lambda}-\boldsymbol{\lambda}_{h})\|=\|Q_{k}\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\leq\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|.$
We now estimate the error of the pressure approximation. For $\boldsymbol{v}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$, by (26),
$(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p-p_{h})$ $=(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p)-(\boldsymbol{f},I_{k,\ell}^{\operatorname{div}}\boldsymbol{v})+(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}_{w}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v},0))$
$=(\operatorname{div}\boldsymbol{\sigma},I_{k,\ell}^{\operatorname{div}}\boldsymbol{v})+(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}_{w}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v},0))$
$=(I_{k}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma},{\rm grad\,}_{h}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}))\!-\!\!\!\sum_{T\in\mathcal{T}_{h}}(\Pi_{F}(I_{k}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma})\boldsymbol{n},\Pi_{F}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}))_{\partial T}$
$\quad+(\boldsymbol{\sigma}_{h}-I_{k}^{\rm tn}\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}_{w}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v},0)).$
Then we get from (7) and (12) that
$(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p-p_{h})\lesssim h^{k+1}|\boldsymbol{u}|_{k+2}|\boldsymbol{v}|_{1}.$
Finally, by the inf-sup condition for $\operatorname{div}$ operator, we have
$\|Q_{\ell}p-p_{h}\|\lesssim\sup_{\boldsymbol{v}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})}\frac{(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p-p_{h})}{|\boldsymbol{v}|_{1}}\lesssim h^{k+1}|\boldsymbol{u}|_{k+2}.$
The estimate (27) demonstrates that the mixed finite element method (24) is pressure-robust.
3.5. $L^{2}$ error estimate
The superconvergence of $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|$ can be obtained through the duality argument. Introduce the dual problem: Find $\tilde{\boldsymbol{u}}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$ and $\tilde{p}\in L^{2}(\Omega)/\mathbb{R}$ satisfy
(28) $\begin{cases}-\Delta\tilde{\boldsymbol{u}}+\nabla\tilde{p}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}&\text{ in }\Omega,\\ \qquad\;\;\operatorname{div}\tilde{\boldsymbol{u}}=0&\text{ in }\Omega.\end{cases}$
We assume that the dual problem (28) has the $H^{2}$-regularity
(29) $\|\tilde{\boldsymbol{u}}\|_{2}\lesssim\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|.$
We refer to [29, Remark I.5.6] and [44, Section 11.5] for the $H^{2}$-regularity (29) in convex domains.
Lemma 3.5.
Let $(\boldsymbol{u},p)$ denote the solution of the Stokes equation (1), and let $(\boldsymbol{u}_{h}$, $\boldsymbol{\lambda}_{h}$, $p_{h})$ be the solution of the mixed finite element method (24). Let $\tilde{\boldsymbol{u}}$ be the solution of the dual problem (28). Assume $\operatorname{skw}{\rm grad\,}\boldsymbol{f}\in L^{2}(\Omega;\mathbb{K})$. We have
(30) $(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}\tilde{\boldsymbol{u}})\lesssim\begin{cases}h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}|\tilde{\boldsymbol{u}}|_{2},&k\geq 1,\\ h^{2}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|\|\tilde{\boldsymbol{u}}\|_{1},&k=0.\end{cases}$
Proof.
It follows from (24a) with $\boldsymbol{v}_{h}=I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}}$ and $\boldsymbol{\mu}_{h}=0$, $\operatorname{div}(I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}})=0$, and (26) that
[…]
That is,
$\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}(I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}}))_{T}-\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h})\boldsymbol{n},[\![\Pi_{F}(I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}})]\!])_{F}=0.$
Then
[…]
$\quad-\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h})\boldsymbol{n},[\![\Pi_{F}(\tilde{\boldsymbol{u}}-I_{k,k}^{\operatorname{div}}\tilde{\boldsymbol{u}})]\!])_{F}.$
When $k\geq 1$, we get from the Cauchy-Schwarz inequality and the estimate (12) of $I_{k,k}^{\operatorname{div}}$ that
$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}\tilde{\boldsymbol{u}})\lesssim h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}|\tilde{\boldsymbol{u}}|_{2}.$
Next consider case $k=0$. Applying the integration by parts to the right hand side of (31) to have
[…]
Since $\operatorname{div}\tilde{\boldsymbol{u}}=0$ and $\operatorname{div}H_{0}^{2}(\Omega;\mathbb{K})=H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ (cf. [24, Theorem 1.1]), there exists a $\tilde{\boldsymbol{\tau}}\in H_{0}^{2}(\Omega;\mathbb{K})$ such that
$\operatorname{div}\tilde{\boldsymbol{\tau}}=\tilde{\boldsymbol{u}},\quad\textrm{and}\quad\|\tilde{\boldsymbol{\tau}}\|_{2}\lesssim\|\tilde{\boldsymbol{u}}\|_{1}.$
Recall the linear finite element space of the differential (d2)(d-2)-form in skew-symmetric form [1, 5]
$\mathbb{V}_{h}^{d-2}:=\{\boldsymbol{\tau}_{h}\in H(\operatorname{div},\Omega;\mathbb{K}):\boldsymbol{\tau}_{h}|_{T}\in\mathbb{P}_{1}(T;\mathbb{K})\;\;\textrm{ for }T\in\mathcal{T}_{h}\}.$
The DoFs for space $\mathbb{V}_{h}^{d-2}$ are given in [13, (3.10)]. Let $I_{h}^{d-2}:H^{2}(\Omega;\mathbb{K})\to\mathbb{V}_{h}^{d-2}$ be the nodal interpolation operator. It holds the following commuting property:
$\operatorname{div}(I_{h}^{d-2}\boldsymbol{\tau})=I_{0,0}^{\operatorname{div}}(\operatorname{div}\boldsymbol{\tau})\quad\forall~\boldsymbol{\tau}\in H^{2}(\Omega;\mathbb{K}).$
Then, employing the integration by parts and (26),
[…]
This together with the interpolation error estimate of $I_{h}^{d-2}$ gives
$(\boldsymbol{\sigma}_{h}-\boldsymbol{\sigma},{\rm grad\,}\tilde{\boldsymbol{u}})\lesssim h^{2}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\||\tilde{\boldsymbol{\tau}}|_{2}\lesssim h^{2}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|\|\tilde{\boldsymbol{u}}\|_{1}.$
This ends the proof. ∎
Theorem 3.6.
Let $(\boldsymbol{u},p)$ denote the solution of the Stokes equation (1), and let $(\boldsymbol{u}_{h}$, $\boldsymbol{\lambda}_{h}$, $p_{h})$ be the solution of the mixed finite element method (24). Assume $\boldsymbol{u}\in H^{k+2}(\Omega;\mathbb{R}^{d})$, $\operatorname{skw}{\rm grad\,}\boldsymbol{f}\in L^{2}(\Omega;\mathbb{K})$, and the $H^{2}$-regularity (29) holds. Then
(32) $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|\lesssim h^{k+2}(|\boldsymbol{u}|_{k+2}+\delta_{k0}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|),$
where $\delta_{00}=1$ and $\delta_{k0}=0$ for $k\geq 1$.
Proof.
Set $\boldsymbol{v}_{h}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$ for ease of presentation. Let $\tilde{\boldsymbol{\sigma}}={\rm grad\,}\tilde{\boldsymbol{u}}$. By the fact $\operatorname{div}\boldsymbol{v}_{h}=0$, and the definition of the operator $I_{k}^{\rm tn}$, we have
$\|\boldsymbol{v}_{h}\|^{2}$ $=-(\operatorname{div}\tilde{\boldsymbol{\sigma}},\boldsymbol{v}_{h})=\sum_{T\in\mathcal{T}_{h}}(\tilde{\boldsymbol{\sigma}},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}-\sum_{T\in\mathcal{T}_{h}}(\Pi_{F}\boldsymbol{v}_{h},\Pi_{F}\tilde{\boldsymbol{\sigma}}\boldsymbol{n})_{\partial T}$
$=\sum_{T\in\mathcal{T}_{h}}(I_{k}^{\rm tn}\tilde{\boldsymbol{\sigma}},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}-\sum_{T\in\mathcal{T}_{h}}(\Pi_{F}\boldsymbol{v}_{h},\Pi_{F}(I_{k}^{\rm tn}\tilde{\boldsymbol{\sigma}})\boldsymbol{n})_{\partial T}$
$=\big(I_{k}^{\rm tn}\tilde{\boldsymbol{\sigma}},\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}\boldsymbol{\lambda}-\boldsymbol{\lambda}_{h})\big).$
This combined with (22) implies
[…]
Then we get from the Cauchy-Schwarz inequality, the estimate (7) of $I_{k}^{\rm tn}$, (30) and the $H^{2}$-regularity (29) that
$\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|\lesssim h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}+\delta_{k0}h^{2}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|.$
Thus, the estimate (32) follows from (27) and the last inequality. ∎
3.6. Postprocessing
We can construct a superconvergent approximation of $\boldsymbol{u}$ using the estimates (27) and (32). For each $T\in\mathcal{T}_{h}$, find $\boldsymbol{u}_{h}^{*}|_{T}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d})$ and $p_{h}^{*}\in\mathbb{P}_{k}(T)\cap L_{0}^{2}(T)$ such that
(33a) $(\boldsymbol{u}_{h}^{*}\cdot\boldsymbol{n},q)_{F}$ $=(\boldsymbol{u}_{h}\cdot\boldsymbol{n},q)_{F}\quad\;\;\;\forall\,q\in\mathbb{P}_{0}(F),\ F\in\mathcal{F}(T),$
(33b) $({\rm grad\,}\boldsymbol{u}_{h}^{*},{\rm grad\,}\boldsymbol{v})_{T}+(\operatorname{div}\boldsymbol{v},p_{h}^{*})_{T}$ $=(\boldsymbol{\sigma}_{h},{\rm grad\,}\boldsymbol{v})_{T}\quad\forall\,\boldsymbol{v}\in\widetilde{\mathbb{P}}_{k+1}(T;\mathbb{R}^{d}),$
(33c) $(\operatorname{div}\boldsymbol{u}_{h}^{*},q)_{T}$ $=0\qquad\qquad\qquad\;\forall\,q\in{\mathbb{P}}_{k}(T)\cap L_{0}^{2}(T),$
where $\widetilde{\mathbb{P}}_{k+1}(T;\mathbb{R}^{d}):=\{\boldsymbol{v}_{h}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d}):Q_{0,F}(\boldsymbol{v}_{h}\cdot\boldsymbol{n})=0\ \ \forall\,F\subset\partial T\}$.
The local problems (33) are well-posed. From (33a) we get
$Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h}^{*})=Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h})=0;$
combined with (33c), this yields $\operatorname{div}\boldsymbol{u}_{h}^{*}=0$ on each $T\in\mathcal{T}_{h}$. Thus $\boldsymbol{u}_{h}^{*}\in\mathbb{P}_{k+1}(\mathcal{T}_{h};\mathbb{R}^{d})$ is pointwise divergence-free elementwise.
Theorem 3.7.
Let $(\boldsymbol{u},p)$ be the solution of the Stokes equation (1), and let $(\boldsymbol{u}_{h}$, $\boldsymbol{\lambda}_{h}$, $p_{h})$ be the solution of the mixed finite element method (24). Assume $\boldsymbol{u}\in H^{k+2}(\Omega;\mathbb{R}^{d})$. Then
(34) $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|\lesssim h^{k+1}|\boldsymbol{u}|_{k+2}.$
If, in addition, $\operatorname{skw}{\rm grad\,}\boldsymbol{f}\in L^{2}(\Omega;\mathbb{K})$ and the $H^{2}$-regularity estimate (29) holds, we have
(35) $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|\lesssim h^{k+2}\big(|\boldsymbol{u}|_{k+2}+\delta_{k0}\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|\big).$
Proof.
Let
$\boldsymbol{w}=(I-I_{(0,0),T}^{\operatorname{div}})(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\in\widetilde{\mathbb{P}}_{k+1}(T;\mathbb{R}^{d}).$
By (15),
$\operatorname{div}\big(I_{(0,0),T}^{\operatorname{div}}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\big)=Q_{0,T}\operatorname{div}\boldsymbol{u}-Q_{0,T}\operatorname{div}\boldsymbol{u}_{h}^{*}=0.$
Hence $I_{(0,0),T}^{\operatorname{div}}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\in\mathbb{P}_{0}(T;\mathbb{R}^{d})$, and applying (15) again gives
$\operatorname{div}\boldsymbol{w}=\operatorname{div}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})=Q_{k,T}\operatorname{div}\boldsymbol{u}-\operatorname{div}\boldsymbol{u}_{h}^{*}=0.$
Using (33b) with $\boldsymbol{v}=\boldsymbol{w}$ yields
$|\boldsymbol{w}|_{1,T}^{2}$ $=({\rm grad\,}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*}),{\rm grad\,}\boldsymbol{w})_{T}$
$=({\rm grad\,}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}),{\rm grad\,}\boldsymbol{w})_{T}+(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}\boldsymbol{w})_{T}.$
Applying the interpolation estimate (12) and the inverse inequality gives
(36) $\|\boldsymbol{w}\|_{T}\eqsim h_{T}|I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*}|_{1,T}\lesssim h_{T}\big(|\boldsymbol{u}-I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}|_{1,T}+\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{T}\big).$
Thus, the estimate (34) follows from the triangle inequality, the interpolation bound (12), and the stress estimate (27).
From (33a),
$I_{(0,0),T}^{\operatorname{div}}(I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}^{*})=I_{(0,0),T}^{\operatorname{div}}(I_{(k,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}).$
Using the triangle inequality, (12), the inverse inequality, and (36), we obtain
$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|_{T}$ $\leq\|\boldsymbol{u}-I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}\|_{T}+\|I_{(0,0),T}^{\operatorname{div}}(I_{(k,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|_{T}+\|\boldsymbol{w}\|_{T}$
$\lesssim\|\boldsymbol{u}-I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}\|_{T}+\|I_{(k,k),T}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|_{T}$
$\quad+h_{T}\big(|\boldsymbol{u}-I_{(k+1,k),T}^{\operatorname{div}}\boldsymbol{u}|_{1,T}+\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{T}\big).$
Combining (12), (32), and (27) gives (35). ∎
4. Equivalent Discrete Methods
In this section, we explore several discrete methods that are equivalent to the mixed finite element scheme (24).
4.1. Hybridization of the velocity
Recall the $H^{1}$-nonconforming virtual element in [7, 15]. For integers $k\geq 0$ and $\ell=k,k-1$, the shape function space is defined as
$V_{k+1,\ell+2}^{\rm VE}(T)=\{v\in H^{1}(T):\Delta v\in\mathbb{P}_{\ell}(T),\partial_{n}v|_{F}\in\mathbb{P}_{k}(F)\textrm{ for }F\in\mathcal{F}(T)\}.$
The DoFs for $V_{k+1,\ell+2}^{\rm VE}(T)$ are given by
(37a) $(v,q)_{F},$ $\quad q\in\mathbb{P}_{k}(F),\,F\in\mathcal{F}(T),$
(37b) $(v,q)_{T},$ $\quad q\in\mathbb{P}_{\ell}(T).$
The global virtual element space $V_{k+1,\ell+2}^{\rm VE}=V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})$ by
[…]
 $\qquad\qquad\qquad\qquad\qquad\quad\quad\;\;\textrm{ single-valued across each face in $\mathring{\mathcal{F}}_{h}$}\}.$
The virtual element space $V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})$ satisfies the weak continuity condition
$([\![v]\!],q)_{F}=0\quad\forall\,v\in V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h}),\ q\in\mathbb{P}_{k}(F),\ F\in\mathring{\mathcal{F}}_{h}.$
When $(k,\ell)=(0,-1)$, the space $V_{1,1}^{\rm VE}$ reduces to the Crouzeix–Raviart (CR) element [25]. For $(k,\ell)=(0,0)$, the space $V_{1,2}^{\rm VE}$ corresponds to the enriched Crouzeix–Raviart element [37, 36].
Find $\boldsymbol{u}_{h}\in\prod_{T\in\mathcal{T}_{h}}\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, $\boldsymbol{\lambda}_{h}\in\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$ and $p_{h}\in V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})/\mathbb{R}$ such that
(38a) $(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h}),\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))-(\boldsymbol{v}_{h},\nabla_{h}p_{h})$ $=(\boldsymbol{f},\boldsymbol{v}_{h}),$
(38b) $(\boldsymbol{u}_{h},\nabla_{h}q_{h})$ $=0$
for all $\boldsymbol{v}_{h}\in\prod_{T\in\mathcal{T}_{h}}\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, $\boldsymbol{\mu}_{h}\in\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$ and $q_{h}\in V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})/\mathbb{R}$.
Theorem 4.1.
The mixed method (38) is well-posed. Let $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},p_{h})\in\prod_{T\in\mathcal{T}_{h}}$ $\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ $\times\mathring{\Lambda}_{k}\times(V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h})/\mathbb{R})$ be its solution, then $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},Q_{\ell}p_{h})\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\times(\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R})$ is the solution of the mixed finite element method (24).
Proof.
We first prove the discrete method (38) has the zero solution when $\boldsymbol{f}=0$. By applying the integration by parts to (38b),
$\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{u}_{h},Q_{\ell,T}q_{h})_{T}-\sum_{F\in\mathcal{F}_{h}}([\![\boldsymbol{u}_{h}\cdot\boldsymbol{n}]\!],Q_{k,F}q_{h})_{F}=0\quad\forall~q_{h}\in V_{k+1,\ell+2}^{\rm VE}(\mathcal{T}_{h}).$
Thanks to DoFs (37), the above equation implies $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ and $\operatorname{div}\boldsymbol{u}_{h}=0$.
Let $Q_{T}^{\rm div}:L^{2}(T;\mathbb{R}^{d})\to\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ be the $L^{2}$-orthogonal projection operator. Recall the following norm equivalences on each $T\in\mathcal{T}_{h}$ established in  [39, 15, 13]:
(39) $\|Q_{T}^{\rm div}\nabla v\|_{0,T}$ $\eqsim\|\nabla v\|_{0,T}\quad\forall\ v\in V_{k+1,\ell+2}^{\rm VE}(T).$
By adding equation (38a) with $(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h})=(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$ and equation (38b) with $q_{h}=p_{h}$, we get $\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})=0$, which means $\boldsymbol{u}_{h}=0$ and $\boldsymbol{\lambda}_{h}=0$. Then we get from equation (38a) and the norm equivalence (39) that $p_{h}=0$. Thus, the discrete method (38) is well-posed.
For the second part, equation (38b) indicates $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ and $\operatorname{div}\boldsymbol{u}_{h}=0$. By restricting $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, it follows from the integration by parts that
$(\boldsymbol{v}_{h},\nabla_{h}p_{h})=(\operatorname{div}\boldsymbol{v}_{h},Q_{\ell}p_{h}).$
Hence, $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},Q_{\ell}p_{h})$ satisfies equation (24a). This ends the proof. ∎
The normal continuity for the velocity space is relaxed in (38) which is particularly useful when $\ell=-1$. As a local basis for the space $\mathring{\mathbb{V}}_{0,-1}^{\operatorname{div}}$ is not readily available, we can implement the equivalent mixed method (38) instead of the mixed method (24).
Even for $(k,\ell)=(0,-1)$, our method has the following estimates:
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}+\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|+\|\operatorname{dev}{\rm grad\,}_{w}(I_{0,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h},Q_{0,\mathcal{F}_{h}}\boldsymbol{\lambda}-\boldsymbol{\lambda}_{h})\|\lesssim h|\boldsymbol{u}|_{2},$
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|_{0}+h\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h})\|\lesssim h(|\boldsymbol{u}|_{2}+|\boldsymbol{u}|_{1}),$
$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|+\|I_{0,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|\lesssim h^{2}(|\boldsymbol{u}|_{2}+\|\operatorname{skw}{\rm grad\,}\boldsymbol{f}\|).$
4.2. Stabilization-free virtual element methods
For $k\geq 0$, define the space of shape functions for an $H^{1}$-nonconforming virtual element as
$\mathbb{V}_{k+1}^{\rm VE}(T):=\{\boldsymbol{v}\in H^{1}(T;\mathbb{R}^{d}):$ $\operatorname{div}\boldsymbol{v}\in\mathbb{P}_{k}(T),\textrm{ there exists some }s\in L^{2}(T)$
$\textrm{ such that }\Delta\boldsymbol{v}+\nabla s\in\big(\mathbb{P}_{k-1}(T;\mathbb{R}^{d})\cap\ker(\cdot\boldsymbol{x})\big),$
$\textrm{ and }(\partial_{n}\boldsymbol{v}+s\boldsymbol{n})|_{F}\in\mathbb{P}_{k}(F;\mathbb{R}^{d})\;\forall~F\in\mathcal{F}(T)\}.$
Following the argument in [53, Section 3.1], the local virtual element space $\mathbb{V}_{k+1}^{\rm VE}(T)$ is uniquely determined by the following DoFs
(40a) $(\boldsymbol{v},\boldsymbol{q})_{F},$ $\quad\boldsymbol{q}\in\mathbb{P}_{k}(F;\mathbb{R}^{d}),F\in\mathcal{F}(T),$
(40b) $(\boldsymbol{v},\boldsymbol{q})_{T},$ $\quad\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d}).$
Clearly we have $\mathbb{P}_{k+1}(T;\mathbb{R}^{d})\subseteq\mathbb{V}_{k+1}^{\rm VE}(T)$, and $\mathbb{V}_{1}^{\rm VE}(T)=\mathbb{P}_{1}(T;\mathbb{R}^{d})$. Then define the global virtual element space by
[…]
A mixed virtual element method for the Stokes equation (1) is to find $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}$ and $p_{h}\in\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$ with integer $k\geq 0$, such that
(41a) $(Q_{k}(\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{u}_{h}),Q_{k}(\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}))+(\operatorname{div}_{h}\boldsymbol{v}_{h},p_{h})$ $=(\boldsymbol{f},I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h}),$
(41b) $(\operatorname{div}_{h}\boldsymbol{u}_{h},q_{h})$ $=0$
for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}$ and $q_{h}\in\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$. When $k=0$, the mixed method (41) is exactly the modified Crouzeix-Raviart element method in [42, (23)].
Lemma 4.2.
It holds
(42) $\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h},Q_{k,\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{v}_{h}))=Q_{k}(\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h})\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}.$
Proof.
By the definitions of the operator $\operatorname{dev}{\rm grad\,}_{w}$ and the operator $I_{k,k}^{\operatorname{div}}$, we have for any $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$ that
$\quad\;(\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h},Q_{k,\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{v}_{h})),\boldsymbol{\tau})_{T}$
$=-(I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot(I_{k,k}^{\operatorname{div}}\boldsymbol{v}_{h}),\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\Pi_{F}\boldsymbol{v}_{h},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}$
$=-(\boldsymbol{v}_{h},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot\boldsymbol{v}_{h},\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\Pi_{F}\boldsymbol{v}_{h},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}.$
Then (42) follows from the integration by parts. ∎
As a consequence of (42), $\|Q_{k}(\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h})\|$ defines a norm on the space $\mathring{\mathbb{V}}_{h}^{\rm VE}\cap\ker(\operatorname{div}_{h})$.
Theorem 4.3.
The mixed virtual element method (41) is well-posed. Let $(\boldsymbol{u}_{h},p_{h})\in\mathring{\mathbb{V}}_{h}^{\rm VE}\times\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$ be its solution. Then $(I_{k,k}^{\operatorname{div}}\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{u}_{h}),Q_{\ell}p_{h})\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h}$; $\mathbb{R}^{d-1})$ $\times\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ be the solution of the mixed finite element method (24).
Proof.
The well-posedness of the mixed finite element method (41) can be proved using an argument analogous to that in Theorem 4.1.
We then prove the equivalence between the discrete method (41) and the discrete method (24). By (42), the equation (41a) means that $(I_{k,k}^{\operatorname{div}}\boldsymbol{u}_{h},Q_{k,\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{u}_{h}),Q_{\ell}p_{h})$ satisfies equation (24a). By the commuting property (15), the equation (41a) indicates that $I_{k,k}^{\operatorname{div}}\boldsymbol{u}_{h}$ satisfies equation (24b). ∎
Remark 4.4.
We can also use the following virtual element space for the velocity:
[…]
4.3. Discrete methods based on the pseudostress-velocity-pressure formulation
By introducing $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\in\Sigma_{k}^{\rm tn}$, we can rewrite the mixed finite element method (24) as the MCS method [31]: find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$, $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ such that
$(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},q_{h};\boldsymbol{u}_{h})$ $=0$ $\forall~\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R},$
$b_{h}(\boldsymbol{\sigma}_{h},p_{h};\boldsymbol{v}_{h})$ $=-(\boldsymbol{f},\boldsymbol{v}_{h})$ $\forall~\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}},$
where the bilinear form
$b_{h}(\boldsymbol{\tau}_{h},q_{h};\boldsymbol{v}_{h})$ $:=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{T}-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n},\boldsymbol{n}\cdot\boldsymbol{v}_{h})_{\partial T}-(\operatorname{div}\boldsymbol{v}_{h},q_{h})$
$\;=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau}_{h},{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{T\in\mathcal{T}_{h}}(\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n},\Pi_{F}\boldsymbol{v}_{h})_{\partial T}-(\operatorname{div}\boldsymbol{v}_{h},q_{h}).$
The subspace
$\left\{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}:\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{d-1})\ \text{for all }F\in\mathcal{F}_{h}\right\}$
was employed in the MCS method [31] to discretize the traceless stress tensor $\boldsymbol{\sigma}$. However, the constraint $\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{d-1})$ prevents the discrete scheme in [31] from achieving superconvergence properties such as the estimates $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|Q_{\ell}p-p_{h}\|$ in (27), and $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|$ in (32).
Superconvergent results were later obtained in [30] by enriching the stress space, but their approach requires $k=\ell\geq 1$, whereas our formulation also covers low-order cases: $(k,\ell)=(0,-1)$, $(0,0)$, and $(1,0)$. Our approach also leads to a conceptually simpler and technically more transparent error analysis.
5. Numerical examples
In this section, we present numerical examples to validate the error estimates of the mixed finite element method (24) and the post-processing procedure (33). The test problems are adapted from [30], which primarily considered higher-order cases with $k=\ell\geq 1$. Here, we focus on low-order settings with $\ell=0$ and $k=0,1$.
The computational domain is $\Omega=(0,1)^{d}$ for $d=2,3$, uniformly partitioned into simplicial meshes. All numerical experiments were developed based on the MATLAB package iiFEM [14].
Example 5.1.
We first test the mixed finite element method (24) in two dimensions. Take the exact solution
$\boldsymbol{u}=\operatorname{curl}\psi_{2},\quad\text{and}\quad p=-x^{5}-y^{5}+\frac{1}{3},$
where $\psi_{2}=x^{2}(x-1)^{2}y^{2}(y-1)^{2}$.
Table 1. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, and $\|p-p_{h}\|$ for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ order $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ order $\|p-p_{h}\|$ order
$2^{-3}$ $(0,0)$ 2.988e-03 3.103e-02 7.810e-02
$2^{-4}$ $(0,0)$ 1.284e-03 1.22 1.677e-02 0.89 3.914e-02 1.00
$2^{-5}$ $(0,0)$ 5.988e-04 1.10 8.700e-03 0.95 1.963e-02 1.00
$2^{-6}$ $(0,0)$ 2.907e-04 1.04 4.440e-03 0.97 9.840e-03 1.00
$2^{-7}$ $(0,0)$ 1.436e-04 1.02 2.247e-03 0.98 4.931e-03 1.00
$2^{-3}$ $(1,0)$ 3.296e-04 2.447e-03 7.453e-02
$2^{-4}$ $(1,0)$ 8.382e-05 1.98 6.305e-04 1.96 3.760e-02 0.99
$2^{-5}$ $(1,0)$ 2.104e-05 1.99 1.597e-04 1.98 1.880e-02 1.00
$2^{-6}$ $(1,0)$ 5.266e-06 2.00 4.016e-05 1.99 9.428e-03 1.00
$2^{-7}$ $(1,0)$ 1.317e-06 2.00 1.007e-05 2.00 4.715e-03 1.00
Table 2. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
$2^{-3}$ $(0,0)$ 1.233e-03 2.890e-02
$2^{-4}$ $(0,0)$ 3.277e-04 1.91 1.481e-02 0.96
$2^{-5}$ $(0,0)$ 8.353e-05 1.97 7.453e-03 0.99
$2^{-6}$ $(0,0)$ 2.099e-05 1.99 3.733e-03 1.00
$2^{-7}$ $(0,0)$ 5.256e-06 2.00 1.867e-03 1.00
$2^{-3}$ $(1,0)$ 3.296e-05 2.286e-03
$2^{-4}$ $(1,0)$ 4.167e-06 2.98 5.183e-04 1.98
$2^{-5}$ $(1,0)$ 5.264e-07 2.99 1.463e-04 1.99
$2^{-6}$ $(1,0)$ 6.625e-08 2.99 3.666e-05 2.00
$2^{-7}$ $(1,0)$ 8.315e-09 2.99 9.178e-06 2.00
Example 5.2.
We then test the mixed finite element method (24) in three dimensions. Take the exact solution
$\boldsymbol{u}=\operatorname{curl}(\psi_{3},\psi_{3},\psi_{3})^{\intercal},\quad\text{and}\quad p=-x^{5}-y^{5}-z^{5}+\frac{1}{2},$
where $\psi_{3}=x^{2}(x-1)^{2}y^{2}(y-1)^{2}z^{2}(z-1)^{2}$.
The numerical errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, and $\|p-p_{h}\|$ are reported in Table 1 for Example 5.1 and Table 3 for Example 5.2, while the post-processed errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ are given in Table 2 for Example 5.1 and Table 4 for Example 5.2. The results in these tables demonstrate that
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|\eqsim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|=\mathcal{O}(h^{k+1}),\quad\|p-p_{h}\|=\mathcal{O}(h),$
and for the post-processed solution $\boldsymbol{u}_{h}^{*}$,
$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|=\mathcal{O}(h^{k+2}),\quad\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|=\mathcal{O}(h^{k+1}),$
which are consistent with the theoretical estimates (27) and (34)–(35).
Table 3. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$, and $\|p-p_{h}\|$ for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ order $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ order $\|p-p_{h}\|$ order
$2^{-1}$ $(0,0)$ 6.762e-04 9.787e-03 2.942e-01
$2^{-2}$ $(0,0)$ 3.110e-04 1.12 4.960e-03 1.06 1.649e-01 0.84
$2^{-3}$ $(0,0)$ 1.399e-04 1.15 2.431e-03 0.95 8.501e-02 0.96
$2^{-4}$ $(0,0)$ 6.719e-05 1.06 1.260e-03 0.95 4.284e-02 0.99
$2^{-5}$ $(0,0)$ 3.320e-05 1.02 6.397e-04 0.98 2.146e-02 1.00
$2^{-1}$ $(1,0)$ 2.076e-04 2.431e-03 2.942e-01
$2^{-2}$ $(1,0)$ 6.700e-05 1.63 5.715e-04 2.09 1.649e-01 0.84
$2^{-3}$ $(1,0)$ 1.790e-05 1.90 1.541e-04 1.89 8.501e-02 0.96
$2^{-4}$ $(1,0)$ 4.551e-06 1.98 3.975e-05 1.96 4.284e-02 0.99
Table 4. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
$2^{-1}$ $(0,0)$ 4.167e-04 4.576e-03
$2^{-2}$ $(0,0)$ 1.565e-04 1.41 2.880e-03 0.67
$2^{-3}$ $(0,0)$ 4.400e-05 1.83 1.539e-03 0.90
$2^{-4}$ $(0,0)$ 1.143e-05 1.95 7.836e-04 0.97
$2^{-5}$ $(0,0)$ 2.889e-06 1.98 3.937e-04 0.99
$2^{-1}$ $(1,0)$ 9.624e-05 1.894e-03
$2^{-2}$ $(1,0)$ 1.493e-05 2.69 5.478e-04 1.79
$2^{-3}$ $(1,0)$ 2.035e-06 2.88 1.486e-04 1.88
$2^{-4}$ $(1,0)$ 2.627e-07 2.95 3.808e-05 1.96

### v2

@@
-Superconvergent and Divergence-Free Finite Element Methods for Stokes Equation
+Superconvergent and Divergence-Free Mixed Finite Element Methods for The Stokes Equation
 Abstract.
-Superconvergent and divergence-free finite element methods for the Stokes equation are developed.
-The velocity and pressure are discretized using $H(\mathrm{div})$-conforming vector elements and discontinuous piecewise polynomials.
-The discrete formulation employs a weak deviatoric gradient operator built with tangential-normal continuous finite elements for traceless tensors, requiring no stabilization.
-Optimal and superconvergent error estimates are established.
-The method connects to nonconforming virtual element and pseudostress-velocity-pressure mixed formulations.
-Numerical experiments verify the theory.
+This paper develops divergence-free mixed finite element methods for the Stokes equation.
+Using H(div)-conforming velocities and discontinuous pressures ensures the inf-sup condition for the velocity–pressure pair and yields pointwise divergence-free velocities.
+However, this choice makes the vector Laplacian difficult to discretize.
+Inspired by mass-conserving mixed formulations with stresses, tangential–normal continuous traceless tensor elements are introduced to discretize the vector Laplacian.
+An inf-sup condition for the weak div operator between the stress and velocity spaces is then proved.
+Two key properties characterize the scheme.
+First, the stress–velocity inf-sup stability gives a stable discretization of the vector Laplacian without additional stabilization, unlike discontinuous Galerkin or virtual element methods.
+Second, the scheme has the property that if a stress field is weakly divergence-free, then it is also strongly divergence-free.
+This decouples the stress and velocity errors and leads to superconvergence.
+As a result, optimal-order error estimates are obtained for the stress, while the velocity and pressure converge at rates higher than the approximation orders of the chosen spaces.
+Numerical experiments confirm the theoretical results.
 Key words and phrases:
@@
 1. Introduction
-In this work, we develop superconvergent and divergence-free finite element methods for the Stokes equation on a bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) with external force $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{d})$:
-(1) $\begin{cases}-\Delta\boldsymbol{u}-\nabla p=\boldsymbol{f}&\text{in }\Omega,\\ \operatorname{div}\boldsymbol{u}=0&\text{in }\Omega,\\ \boldsymbol{u}=0&\text{on }\partial\Omega,\end{cases}$
-where $\boldsymbol{u}$ and $p$ denote the velocity and pressure.
-Classical stable finite element pairs for (1) include the Taylor–Hood element [50], the MINI element [3], and the nonconforming $P_{1}$$P_{0}$ element [25].
-More examples and a detailed discussion are given in [10].
-However, these classical pairs are not pointwise divergence-free, which weakens mass conservation and reduces pressure robustness in error estimates.
-A review of the importance of divergence-free constraints and their numerical implications can be found in [41].
-Divergence-free finite element pairs include the Scott–Vogelius element [48], which requires special mesh conditions; smooth finite element pairs [56, 17, 18] with supersmooth degrees of freedom; conforming pairs on split meshes [6, 55, 22, 28, 32, 38]; conforming rational pairs [33, 34]; and low-order, normal-continuous nonconforming pairs [43, 49, 54].
-We also refer to [8, 52, 51] for discontinuous Galerkin methods, and to [9, 20, 53, 40] for virtual element methods.
-One can impose the divergence-free condition by using an $H(\operatorname{div})$-conforming finite element space for the velocity.
-Specifically, we use the Raviart–Thomas (RT) element space [47, 45] or the Brezzi–Douglas–Marini (BDM) element space [12, 46]:
-$\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}:=\{\boldsymbol{v}_{h}\in H_{0}(\operatorname{div},\Omega):\boldsymbol{v}_{h}|_{T}\in\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x}\;\text{ for }T\in\mathcal{T}_{h}\}$
-for the velocity, with integers $\ell=k$ (${\rm RT}_{k}$) or $\ell=k-1$ (${\rm BDM}_{k}$), and the discontinuous piecewise polynomial space $\mathbb{P}_{\ell}(\mathcal{T}_{h})$ for the pressure.
-Since $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\not\subset H^{1}(\Omega;\mathbb{R}^{d})$, the discretization of $\Delta\boldsymbol{u}$ is nontrivial.
-One approach is to use discontinuous Galerkin (DG) methods for the Laplacian operator, as in [52, 23].
-The DG method in [52] introduces a penalty term to ensure stability, while the hybridizable DG (HDG) method in [23] employs $k$th-order polynomial spaces for all variables.
-In fact, it suffices to take $p_{h}\in\mathbb{P}_{k-1}(\mathcal{T}_{h})$ when the velocity $\boldsymbol{u}_{h}\in{\rm BDM}_{k}$.
-In the vorticity–velocity–pressure formulation [26, 27], the vorticity $\omega={\rm curl\,}\boldsymbol{u}$ is introduced, and
-$\Delta\boldsymbol{u}={\rm curl\,}\omega-{\rm grad\,}\operatorname{div}\boldsymbol{u}={\rm curl\,}{\rm curl\,}\boldsymbol{u}-{\rm grad\,}\operatorname{div}\boldsymbol{u}$
-is interpreted as the Hodge Laplacian.
-However, as observed in [26, 27], for no-slip Dirichlet boundary conditions ($\boldsymbol{u}=0$ on $\partial\Omega$), there are stability issues and reduced convergence rates.
-This was later explained in [4]: the suboptimal convergence rate comes from the lack of a proper Hilbert complex in this setting.
-The lost convergence rate can be partly recovered by mesh symmetry.
-In the triangular MAC (TMAC) schemes for the Stokes equation [21], mass lumping is applied to the vorticity so that it can be eliminated, and the velocity and pressure convergence rates are improved by mesh symmetry.
-Using the element pair $\textrm{RT}_{0}$$\textrm{P}_{0}$ for velocity and pressure (i.e., $k=\ell=0$) yields
-$|I_{0,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}|_{1,h}+\|Q_{0}p-p_{h}\|\lesssim h^{\min\{1,s\}}|\ln h|^{1/2},$
-where $s\geq\tfrac{1}{2}$ measures mesh symmetry and $I^{\rm\operatorname{div}}_{k,\ell}$ is the interpolation operator into $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$.
-With the pair $\textrm{BDM}_{1}$$\textrm{P}_{0}$ (i.e., $k=1$, $\ell=0$), the TMAC scheme achieves
-$|I_{1,0}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}|_{1,h}+\|Q_{0}p-p_{h}\|\lesssim h^{1+\min\{1/2,s\}}.$
-In [30, 31], the mass-conserving mixed formulation with stresses (MCS) finite element method was developed for the pseudostress–velocity–pressure formulation, also employing $H(\operatorname{div})$-conforming velocity spaces.
-The stress space in [31] is slightly nonconforming, leading to the standard $h^{k}$ convergence instead of the superconvergent $h^{k+1}$ rate.
-In [30], a weak symmetry constraint was introduced and the stress space was enriched, yielding $h^{k+1}$ convergence for the $\mathrm{RT}_{k}$$\mathrm{P}_{k}$ pair with $k\geq 1$.
-A divergence free weak virtual element method was developed in [20], which requires stabilization and does not yield superconvergence.
-The main idea is to introduce a piecewise vector-valued polynomial $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}:=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$ defined on faces to enforce tangential continuity.
-We follow this approach and improve it to a stabilization-free scheme with superconvergence.
-Define the weak deviatoric gradient operator element-wise by
-$(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v},\boldsymbol{\mu}),\boldsymbol{\tau})_{T}=-(\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}+(\boldsymbol{n}\cdot\boldsymbol{v},\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}+(\boldsymbol{\mu},\Pi_{F}\boldsymbol{\tau}\boldsymbol{n})_{\partial T}$
-for any $\boldsymbol{\tau}\in\mathbb{P}_{k}(T;\mathbb{T})$ and $T\in\mathcal{T}_{h}$.
-The tensor space is reduced to the traceless tensor $\mathbb{T}$, since for $\boldsymbol{\sigma}:={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{tr}{\rm grad\,}\boldsymbol{u}=\operatorname{div}\boldsymbol{u}=0$.
-We propose the following mixed finite element method: find $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\lambda}_{h}\in\mathring{\Lambda}_{k}$, and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ with integers $k\geq 0$ and $\ell\in\{k,k-1\}$, such that
-(2a) $(\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h}),\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))+(\operatorname{div}\boldsymbol{v}_{h},p_{h})$ $=(\boldsymbol{f},\boldsymbol{v}_{h}),$
-(2b) $(\operatorname{div}\boldsymbol{u}_{h},q_{h})$ $=0$
-for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, $\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}$, and $q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$.
-Notably, the discrete method (2) requires no additional stabilization, despite relying on the weak differential operator $\operatorname{dev}{\rm grad\,}_{w}$.
-We derive superconvergent estimates for the velocity and pressure:
-$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|\operatorname{dev}{\rm grad\,}_{w}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h})\|+\|Q_{\ell}p-p_{h}\|\lesssim h^{k+1}|\boldsymbol{\sigma}|_{k+1},$
-where $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u},$ and $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h})$.
-Unlike the TMAC scheme [21], such superconvergence does not require the symmetry of the triangulation.
-It is further utilized to construct a new superconvergent approximation of $\boldsymbol{u}$.
-Our method allows $\ell=0$ an important case not covered by MCS [30, 31].
-We provide an illustration of the case $(k,\ell)=(0,0)$, which can be treated as a generalization of the classical MAC scheme [35] on rectangular grids to triangulations and achieve first order $h$ convergence, no loss of convergence compared with the vorticity-velocity-pressure formulation.
-One can use $\textrm{BDM}_{1}$$\textrm{P}_{0}$ pair, $(k,\ell)=(1,0)$, to get a second order $h^{2}$ scheme.
-We propose a hybridization of (2) by relaxing the normal continuity of the velocity, which enables the choice $(k,\ell)=(0,-1)$.
-The velocity is approximated by the piecewise constant vector space, and the normal continuity is enforced through a nonconforming linear element space for the pressure $p_{h}$.
-Even in this low-order setting, the method achieves at least first-order convergence, which is a notable feature.
-The remainder of this paper is organized as follows.
-Section 2 discusses the weak div stability and the distributional discretization of the vector Laplacian.
-Section 3 develops mixed finite element methods for the Stokes equation, and Section 4 presents their equivalent discrete formulations.
-Finally, Section 5 reports numerical experiments that confirm and illustrate the theoretical results.
+In this work, we develop mixed finite element methods for the Stokes equations on a bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) with a given body force $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{d})$:
+(1) $-\Delta\boldsymbol{u}+\nabla p$ $=\boldsymbol{f}$ $\text{in }\Omega,$
+$\operatorname{div}\boldsymbol{u}$ $=0$ $\text{in }\Omega,$
+$\boldsymbol{u}$ $=\boldsymbol{0}$ $\text{on }\partial\Omega,$
+where $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$ and $p\in L^{2}_{0}(\Omega)$ represent the velocity and pressure, respectively.
+Standard velocity–pressure finite element schemes for the Stokes equations depend crucially on the discrete divergence operator induced by the bilinear form $(\operatorname{div}\boldsymbol{v}_{h},q_{h})$.
+A stable scheme requires the discrete spaces $\boldsymbol{V}_{h}$ and $P_{h}$ to satisfy the inf-sup (or LBB) condition [12]:
+(2) $\inf_{q_{h}\in P_{h}\setminus\{0\}}\sup_{\boldsymbol{v}_{h}\in\boldsymbol{V}_{h}\setminus\{\boldsymbol{0}\}}\frac{(\operatorname{div}\boldsymbol{v}_{h},q_{h})}{\|\boldsymbol{v}_{h}\|_{\boldsymbol{V}_{h}}\|q_{h}\|_{P_{h}}}\geq\beta>0,$
+where $\beta$ is a constant independent of the mesh size.
+Standard stable pairs for (1) include the Taylor–Hood element [52], the MINI element [3], and the nonconforming $P_{1}$$P_{0}$ element [26]; a comprehensive review is available in [10].
+However, these elements do not enforce the divergence-free condition pointwise.
+Losing exact mass conservation often leads to poor pressure robustness in the velocity error estimates.
+We refer to [42] for a detailed discussion on the importance of strong divergence-free constraints.
+Enforcing the discrete divergence-free condition, $(\operatorname{div}\boldsymbol{u}_{h},q_{h})=0$ for all $q_{h}\in P_{h}$, does not yield a strong divergence-free velocity field (i.e., $\operatorname{div}\boldsymbol{u}_{h}=0$ pointwise) because $P_{h}$ is only a proper subspace of $L^{2}_{0}(\Omega)$.
+Schemes that achieve this pointwise property are called strong divergence-free.
+In such schemes, one can choose an appropriate interpolant $\boldsymbol{u}_{I}$ such that $\operatorname{div}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h})=0$ to decouple the velocity error from the pressure error via Galerkin orthogonality:
+(3) $(\nabla(\boldsymbol{u}-\boldsymbol{u}_{h}),\nabla(\boldsymbol{u}_{I}-\boldsymbol{u}_{h}))=(p-p_{h},\operatorname{div}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h}))=0.$
+If the scheme is not strong divergence-free, the term $(p,\operatorname{div}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h}))$ does not generally vanish, introducing a pressure dependency into the velocity error.
+When it does vanish, we obtain the bound $\|\nabla(\boldsymbol{u}-\boldsymbol{u}_{h})\|\leq\|\nabla(\boldsymbol{u}-\boldsymbol{u}_{I})\|$, which depends only on the regularity of the velocity.
+This decoupling yields a pressure-robust error estimate.
+Several methods provide strong divergence-free velocity approximations.
+These include the Scott–Vogelius element [50] (which requires specific mesh conditions), smooth finite elements [57, 18, 19], conforming pairs on split meshes [7, 56, 23, 29, 33, 39], and conforming rational functions [34, 35].
+Low-order nonconforming pairs [45, 51, 55] are also used.
+A natural way to enforce the strong divergence-free condition is to employ $H(\operatorname{div})$-conforming velocity spaces together with discontinuous polynomial pressure spaces.
+For example, using the Raviart–Thomas ($\mathrm{RT}_{k}$) [49, 47] or Brezzi–Douglas–Marini ($\mathrm{BDM}_{k}$) [13, 48] spaces for velocity (with $\boldsymbol{u}\cdot\boldsymbol{n}=0$ on $\partial\Omega$) paired with the discontinuous space $\mathbb{P}_{\ell}(\mathcal{T}_{h})$ for pressure ensures LBB stability and yields a strongly divergence-free velocity field.
+However, because the discrete velocity space is not a subspace of $H^{1}(\Omega;\mathbb{R}^{d})$, the vector Laplacian $\Delta\boldsymbol{u}$ lacks a standard $H^{1}$ weak formulation.
+To address this, the vorticity–velocity–pressure formulation [27, 28] introduces the vorticity $\boldsymbol{\omega}=\operatorname{curl}\boldsymbol{u}$ and treats the Laplacian as the Hodge Laplacian: $\Delta\boldsymbol{u}=\operatorname{curl}\operatorname{curl}\boldsymbol{u}-\nabla\operatorname{div}\boldsymbol{u}$.
+Yet, when enforcing no-slip boundary conditions ($\boldsymbol{u}=\boldsymbol{0}$), this formulation suffers from reduced stability and suboptimal convergence rates due to the absence of a proper Hilbert complex [4].
+Symmetrical meshes can partially recover these rates, as seen in triangular MAC (TMAC) schemes [22], which lump mass to eliminate vorticity.
+Alternatively, the mass-conserving mixed stress (MCS) formulation [31, 32] embeds $H(\operatorname{div})$ velocities in a pseudostress–velocity–pressure system.
+The finite element space for the stress is tangential–normal continuous.
+While standard stress formulations yield $\mathcal{O}(h^{k})$ convergence [32], enriching the stress space allows the $\mathrm{RT}_{k}$$\mathrm{P}_{k}$ pair to achieve $\mathcal{O}(h^{k+1})$ convergence, with requirement $k\geq 1$ [31].
+A sophisticated inf-sup condition is proved for tangential–normal continuous traceless tensors.
+Inspired by the MCS framework, we use tangential–normal continuous elements, $\Sigma_{k}^{\operatorname{tn}}$, to discretize the vector Laplacian.
+Because $\Sigma_{k}^{\operatorname{tn}}$ is not $H(\operatorname{div})$-conforming, we define a weak divergence operator, $\operatorname{div}_{w}$, whose adjoint is the weak deviatoric gradient, $\operatorname{dev}\operatorname{grad}_{w}$, acting on the velocity element.
+The viability of this method depends on establishing an inf-sup condition for $\operatorname{div}_{w}$ over the stress–velocity pair.
+This presents a structural tension: the velocity space must be large enough (e.g., $\mathrm{RT}_{k}$) to maintain the strong divergence-free property—which is necessary to decouple the stress and velocity approximation errors—but small enough (e.g., $\mathrm{BDM}_{k}$) to ensure that $\operatorname{div}_{w}\Sigma_{k}^{\operatorname{tn}}$ remains surjective.
+We resolve this conflict by establishing the inf-sup condition directly on the divergence-free subspace:
+(4) $\inf_{\boldsymbol{v}_{h}\in\mathrm{BDM}_{k}\cap\ker(\operatorname{div})}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\operatorname{tn}}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|\boldsymbol{v}_{h}\|}=\alpha>0,$
+leveraging the crucial algebraic identity $\mathrm{BDM}_{k}\cap\ker(\operatorname{div})=\mathrm{RT}_{k}\cap\ker(\operatorname{div})$.
+Space $\mathrm{RT}_{k}$ is an enrichment of $\mathrm{BDM}_{k}$ to enlarge the range of the divergence operator.
+Furthermore, we establish the strong divergence-free property: if $(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}=0$ for all $\boldsymbol{v}_{h}\in\mathrm{BDM}_{k}\cap\ker(\operatorname{div})$, then in the distributional sense:
+$(\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v})=0\qquad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div}).$
+The constraint $\operatorname{div}\boldsymbol{u}_{h}=0$ is imposed by introducing a Lagrange multiplier, which has the physical interpretation of pressure.
+Specifically, let $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ denote the $\mathrm{RT}_{k}$ space when $\ell=k$, or the $\mathrm{BDM}_{k}$ space when $\ell=k-1$, satisfying the boundary condition $\boldsymbol{u}\cdot\boldsymbol{n}=0$ on $\partial\Omega$.
+The discrete problem seeks $(\boldsymbol{u}_{h},p_{h})\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathbb{P}_{\ell}(\mathcal{T}_{h})$ for $k\geq 0$ and $\ell\in\{k,k-1\}$ such that:
+(5) $\begin{cases}a_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h})+b(\boldsymbol{v}_{h},p_{h})=(\boldsymbol{f},\boldsymbol{v}_{h})&\text{for all }\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}},\\ b(\boldsymbol{u}_{h},q_{h})=0&\text{for all }q_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h}),\end{cases}$
+where $a_{h}(\boldsymbol{u}_{h},\boldsymbol{v}_{h}):=(\operatorname{dev}\operatorname{grad}_{w}\boldsymbol{u}_{h},\operatorname{dev}\operatorname{grad}_{w}\boldsymbol{v}_{h})$ and $b(\boldsymbol{v}_{h},q_{h}):=(\operatorname{div}\boldsymbol{v}_{h},q_{h})$.
+Section 2 details the precise definitions of these spaces and the $\operatorname{dev}\operatorname{grad}_{w}$ operator, while Section 3 presents a hybridized version of (5).
+The boundary condition of the tangential component $\boldsymbol{u}\cdot\boldsymbol{t}=0$ on $\partial\Omega$ is imposed weakly and can be replaced by $\boldsymbol{t}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}=0$ on $\partial\Omega$.
+We emphasize that this method requires no additional stabilization.
+The variational formulation is not augmented by mesh-dependent penalty parameters or grad-div stabilization terms.
+Instead, stability follows directly from the discrete inf-sup conditions for the $\operatorname{div}_{w}$ (4) and $\operatorname{div}$ operators (2).
+The weak deviatoric gradient, $\operatorname{dev}\operatorname{grad}_{w}$, naturally incorporates face contributions via elementwise integration by parts, yielding a consistent, parameter-free discretization that recovers the continuous identity for smooth functions.
+Let $\boldsymbol{\sigma}=\operatorname{dev}\operatorname{grad}\boldsymbol{u}$ and $\boldsymbol{\sigma}_{h}=\operatorname{dev}\operatorname{grad}_{w}\boldsymbol{u}_{h}\in\Sigma_{k}^{\operatorname{tn}}$ denote the continuous and discrete stresses, respectively.
+Owing to the strong divergence-free property of $\operatorname{div}_{w}$, the stress error decouples from the velocity and pressure (using argument similar to (3)), achieving optimal $\mathcal{O}(h^{k+1})$ convergence in $L^{2}$.
+This yields the following estimate:
+$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|+\|\operatorname{dev}\operatorname{grad}_{w}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h})\|+\|Q_{\ell}p-p_{h}\|\lesssim h^{k+1}|\boldsymbol{\sigma}|_{k+1},$
+where $\boldsymbol{u}_{I}$ is an interpolant of $\boldsymbol{u}$ into $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$, and $Q_{\ell}$ is the $L^{2}$-orthogonal projection onto $\mathbb{P}_{\ell}(\mathcal{T}_{h})$.
+Because $h^{k+1}$ exceeds the standard approximation capabilities of the chosen velocity and pressure spaces, both the velocity error $\|\operatorname{dev}\operatorname{grad}_{w}(\boldsymbol{u}_{I}-\boldsymbol{u}_{h})\|$ and the pressure error $\|Q_{\ell}p-p_{h}\|$ are superconvergent.
+This structural superconvergence does not require mesh symmetry—unlike TMAC [22]—and it facilitates higher-order velocity post-processing.
+Alternatively, discontinuous Galerkin (DG) methods can discretize the vector Laplacian [53, 43, 24].
+However, the DG method in [53] requires a penalty term for stability, and the hybridizable DG (HDG) method in [24] uses $k$th-order polynomials for all variables, whereas for $\boldsymbol{u}_{h}\in\mathrm{BDM}_{k}$, it suffices to choose $p_{h}\in\mathbb{P}_{k-1}(\mathcal{T}_{h})$.
+Similarly, the divergence-free weak virtual element method (VEM) [21, 9, 54, 41] generally requires stabilization for the projected discrete gradient and relies on the full $\mathrm{BDM}_{k}$ space; using the incomplete $\mathrm{RT}_{k}$ space degrades convergence.
+Consequently, their energy-norm velocity estimates are optimal, but not superconvergent.
+Our framework extends to the lowest-order case $\ell=0$, which the MCS framework does not address [31, 32].
+Specifically, the $(k,\ell)=(0,0)$ configuration (illustrated in Fig.
+1) generalizes the classical MAC scheme [36] to unstructured triangulations.
+It achieves the first-order convergence without the accuracy loss typical of vorticity-based formulations.
+Similarly, the $\mathrm{BDM}_{1}$$\mathbb{P}_{0}$ pair ($(k,\ell)=(1,0)$) provides a second-order scheme.
+Figure 1.
+Lowest-order configuration $(k,\ell)=(0,0)$.
+The velocity–pressure pair is ${\rm RT}_{0}$$\mathbb{P}_{0}$, and the stress is a piecewise constant traceless matrix with tangential–normal continuity.
+We can further hybridize (5) by relaxing the normal continuity of the velocity, which permits the choice $(k,\ell)=(0,-1)$.
+Here, the velocity is approximated by piecewise constant vectors, and normal continuity is enforced via a nonconforming linear pressure space.
+This configuration maintains the first-order convergence.
+[…]
+Although $H(\operatorname{div})$-conforming velocity fields naturally enforce mass conservation, the primary challenge remains the consistent discretization of the vector Laplacian.
+We address this by developing a distributional framework in Section 2 and establishing the inf-sup condition for the stress–velocity pairs.
+In Section 3, we formulate the mixed finite element methods for the Stokes equations by introducing the pressure as a Lagrange multiplier to enforce the divergence-free constraint.
+We prove the stability of $\operatorname{div}_{w}$ onto $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, including the divergence-free subspace $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$, and establish optimal-order error estimates.
+Furthermore, Section 4 explores equivalent discrete formulations, including virtual element and pseudostress methods.
+Finally, Section 5 provides numerical experiments to verify our theoretical findings, and Section 6 offers concluding remarks.
 2. Distributional Discretization of the Vector Laplacian
-In this section, we discretize the weak divergence operator
-$\operatorname{div}:L^{2}(\Omega;\mathbb{T})\to H^{-1}(\Omega;\mathbb{R}^{d}),$
-and apply it to the vector Laplacian problem.
-We introduce weak divergence and deviatoric gradient operators, establish their inf-sup stability, and derive a superconvergent error estimate for divergence-free velocity spaces.
+The Stokes equation can be viewed as a vector Poisson problem on the divergence-free subspace, i.e., $\Delta\boldsymbol{u}=\boldsymbol{f}$ for $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div}).$ However, $H(\operatorname{div})$ velocity fields do not have the $H^{1}$ regularity needed for a classical Galerkin treatment of the Laplacian.
+This section develops the distributional framework needed to address this issue.
+By rewriting the Laplacian as a first-order system with a traceless tensor (pseudostress), weak divergence and weak deviatoric gradient operators, $\mathrm{div}_{w}$ and $\operatorname{dev}{\rm grad\,}_{w}$, are introduced.
+Their stability and commuting properties are established to ensure consistency with the underlying integration-by-parts identities.
+With this mixed approximation of the vector Laplacian in place, the full Stokes system in Section 3 follows by enforcing the divergence-free constraint through a pressure Lagrange multiplier.
 2.1. Notation
 Let $\Omega\subset\mathbb{R}^{d}$ ($d\geq 2$) be a bounded polytope with boundary $\partial\Omega$.
-For a bounded domain $D$ and integer $m\geq 0$, let $H^{m}(D)$ be the Sobolev space with norm $\|\cdot\|_{m,D}$ and semi-norm $|\cdot|_{m,D}$, and let $H_{0}^{m}(D)$ be the closure of $C_{0}^{\infty}(D)$ under $\|\cdot\|_{m,D}$.
-Set $L^{2}(D)=H^{0}(D)$ with inner product $(\cdot,\cdot)_{D}$, and write $\|\cdot\|_{0,D}$ as $\|\cdot\|_{D}$.
+For a bounded domain $D$ and an integer $m\geq 0$, we denote by $H^{m}(D)$ the standard Sobolev space with norm $\|\cdot\|_{m,D}$ and semi-norm $|\cdot|_{m,D}$.
+Let $H_{0}^{m}(D)$ be the closure of $C_{0}^{\infty}(D)$ in $H^{m}(D)$.
+We set $L^{2}(D)=H^{0}(D)$ with inner product $(\cdot,\cdot)_{D}$ and norm $\|\cdot\|_{D}$.
 When $D=\Omega$, the subscript $D$ is omitted.
-Let $h_{D}$ denote the diameter of $D$, and $\boldsymbol{n}_{\partial D}$ the unit outward normal, abbreviated as $\boldsymbol{n}$.
-We use
+For any $D$, $h_{D}$ denotes its diameter and $\boldsymbol{n}_{\partial D}$ its unit outward normal vector.
+It will be abbreviated as $\boldsymbol{n}$, when the domain $D$ is clear in the context.
+We define the following spaces for $H(\operatorname{div})$:
 $H(\operatorname{div},D)$ $=\{\boldsymbol{v}\in L^{2}(D;\mathbb{R}^{d}):\operatorname{div}\boldsymbol{v}\in L^{2}(D)\},$
-$H_{0}(\operatorname{div},D)$ $=\{\boldsymbol{v}\in H(\operatorname{div},D):\boldsymbol{v}\cdot\boldsymbol{n}=0\text{ on }\partial D\}.$
-Let $L_{0}^{2}(D)=\{q\in L^{2}(D):\int_{D}q\,\,{\rm d}x=0\}.$ For any $B\subseteq L^{2}(D)$, denote $B/\mathbb{R}=B\cap L_{0}^{2}(D)$.
-Let $\{\mathcal{T}_{h}\}_{h>0}$ be a regular family of simplicial meshes of $\Omega$, where $h=\max_{T\in\mathcal{T}_{h}}h_{T}$ and $h_{T}=\operatorname{diam}(T)$.
-Denote by $\mathcal{F}_{h}$ and $\mathring{\mathcal{F}}_{h}$ the sets of all faces and interior faces, respectively.
-For each face $F\in\mathcal{F}_{h}$, fix a unit normal $\boldsymbol{n}_{F}$.
-For $T\in\mathcal{T}_{h}$, let $\mathcal{F}(T)$ be the set of all (d1)(d-1)-dimensional faces of $T$.
-For a face $F$ and vector $\boldsymbol{v}\in\mathbb{R}^{d}$, define its tangential projection
+$H_{0}(\operatorname{div},D)$ $=\{\boldsymbol{v}\in H(\operatorname{div},D):\boldsymbol{v}\cdot\boldsymbol{n}=0\text{ on }\partial D\},$
+endowed with the norm $\|\boldsymbol{v}\|_{\operatorname{div}}=(\|\boldsymbol{v}\|^{2}+\|\operatorname{div}\boldsymbol{v}\|^{2})^{1/2}$.
+Let $L_{0}^{2}(D)=\{q\in L^{2}(D):\int_{D}q\,\,{\rm d}x=0\}$.
+For any subspace $V\subseteq L^{2}(D)$, we denote $V/\mathbb{R}=V\cap L_{0}^{2}(D)$.
+Let $\{\mathcal{T}_{h}\}_{h>0}$ be a regular family of simplicial meshes of $\Omega$, with $h=\max_{T\in\mathcal{T}_{h}}h_{T}$ and $h_{T}=\operatorname{diam}(T)$.
+We denote by $\mathcal{F}_{h}$ and $\mathring{\mathcal{F}}_{h}$ the sets of all faces and interior faces, respectively.
+For any $T\in\mathcal{T}_{h}$, $\mathcal{F}(T)$ denotes the set of faces of $T$.
+For each interior face $F=T^{+}\cap T^{-}$, we fix a unit normal $\boldsymbol{n}_{F}$ to orient jumps and averages.
+Recall that $\boldsymbol{n}_{\partial T}$ is the outward unit normal on the boundary of an element $T$.
+On an interior face $F$, we set $\boldsymbol{n}_{F}=\boldsymbol{n}_{\partial T^{+}}|_{F}$ and thus $\boldsymbol{n}_{\partial T^{-}}|_{F}=-\boldsymbol{n}_{F}$.
+For boundary faces $F\subset\partial\Omega$, we set $\boldsymbol{n}_{F}=\boldsymbol{n}_{\partial\Omega}|_{F}$.
+For two adjacent elements $T^{\pm}$ sharing a face $F$, the jump of a function $v$ is defined as
+$\llbracket v\rrbracket|_{F}=v^{+}\boldsymbol{n}_{F}\cdot\boldsymbol{n}_{\partial T^{+}}+v^{-}\boldsymbol{n}_{F}\cdot\boldsymbol{n}_{\partial T^{-}},$
+and on boundary faces $F\subset\partial\Omega$, we set $\llbracket v\rrbracket=v|_{F}$.
+For a face $F$ and a vector $\boldsymbol{v}\in\mathbb{R}^{d}$, the tangential projection is
 $\Pi_{F}\boldsymbol{v}=(\boldsymbol{I}-\boldsymbol{n}_{F}\boldsymbol{n}_{F}^{\intercal})\boldsymbol{v}.$
-For two adjacent elements $T^{\pm}$ sharing $F$, define the jump
-$\llbracket v\rrbracket|_{F}=v^{+}\boldsymbol{n}_{F}\cdot\boldsymbol{n}_{\partial T^{+}}+v^{-}\boldsymbol{n}_{F}\cdot\boldsymbol{n}_{\partial T^{-}},$
-and on a boundary face $F\subset\partial\Omega$, set $\llbracket v\rrbracket=v|_{F}\,\boldsymbol{n}_{F}\cdot\boldsymbol{n}$.
 For a domain $D$ and integer $k\geq 0$, let $\mathbb{P}_{k}(D)$ be the space of polynomials on $D$ of degree $\leq k$, $\mathbb{H}_{k}(D)=\mathbb{P}_{k}(D)\setminus\mathbb{P}_{k-1}(D)$, and $Q_{k,D}$ the $L^{2}$-projection onto $\mathbb{P}_{k}(D)$.
-Set $\mathbb{M}=\mathbb{R}^{d\times d}$, $\mathbb{K}$ for skew-symmetric, and $\mathbb{T}$ for traceless matrices.
-For $\boldsymbol{\tau}\in\mathbb{M}$,
-$\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\tfrac{1}{d}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I},\qquad\operatorname{skw}\boldsymbol{\tau}=\tfrac{1}{2}(\boldsymbol{\tau}-\boldsymbol{\tau}^{\intercal}).$
+We understand $\mathbb{P}_{-1}(D)=\{0\}$.
+Set $\mathbb{M}=\mathbb{R}^{d\times d}$, and define $\mathbb{K}$ and $\mathbb{T}$ to be the skew-symmetric and traceless subspaces of $\mathbb{M}$, respectively.
+For $\boldsymbol{\tau}\in\mathbb{M}$, define
+$\operatorname{dev}\boldsymbol{\tau}=\boldsymbol{\tau}-\tfrac{1}{d}(\operatorname{tr}\boldsymbol{\tau})\boldsymbol{I}\in\mathbb{T},\qquad\operatorname{skw}\boldsymbol{\tau}=\tfrac{1}{2}\bigl(\boldsymbol{\tau}-\boldsymbol{\tau}^{\intercal}\bigr)\in\mathbb{K}.$
 Define
-$H^{s}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in H^{s}(T)\}=\prod_{T\in\mathcal{T}_{h}}H^{s}(T),$
-$\mathbb{P}_{k}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in\mathbb{P}_{k}(T)\}=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T),$
-$\mathbb{P}_{k}(\mathcal{F}_{h})=\{v\in L^{2}(\mathcal{F}_{h}):v|_{F}\in\mathbb{P}_{k}(F)\}=\prod_{F\in\mathcal{F}_{h}}\mathbb{P}_{k}(F),$
-and similarly define $\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$ for interior faces only. For a space $B(D)$ or $B(\mathcal{T}_{h})$, define
-$B(D;\mathbb{X})=B(D)\otimes\mathbb{X},\quad B(\mathcal{T}_{h};\mathbb{X})=B(\mathcal{T}_{h})\otimes\mathbb{X},$
+$H^{s}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in H^{s}(T)\;\textrm{ for all }T\in\mathcal{T}_{h}\}=\prod_{T\in\mathcal{T}_{h}}H^{s}(T),$
+$\mathbb{P}_{k}(\mathcal{T}_{h})=\{v\in L^{2}(\Omega):v|_{T}\in\mathbb{P}_{k}(T)\;\textrm{ for all }T\in\mathcal{T}_{h}\}=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T),$
+$\mathbb{P}_{k}(\mathcal{F}_{h})=\{v\in L^{2}(\mathcal{F}_{h}):v|_{F}\in\mathbb{P}_{k}(F)\;\textrm{ for all }F\in\mathcal{F}_{h}\}=\prod_{F\in\mathcal{F}_{h}}\mathbb{P}_{k}(F),$
+and similarly define $\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$ for interior faces only. For a linear space $V(D)$ or $V(\mathcal{T}_{h})$, we denote its vector- or tensor-valued version by
+$V(D;\mathbb{X})=V(D)\otimes\mathbb{X},\quad V(\mathcal{T}_{h};\mathbb{X})=V(\mathcal{T}_{h})\otimes\mathbb{X},$
 where $\mathbb{X}\in\{\mathbb{R}^{d},\mathbb{M},\mathbb{T},\mathbb{K}\}$.
-Let $Q_{k}$ denote the $L^{2}$-projection operator from $L^{2}(\Omega)$ onto $\mathbb{P}_{k}(\mathcal{T}_{h})$; its vector- or tensor-valued version is also denoted by $Q_{k}$.
-We use ${\rm grad\,}_{h}$ ($\nabla_{h}$ for scalar function) and $\operatorname{div}_{h}$ to represent the element-wise gradient and divergence operators, respectively.
+Let $Q_{k}$ be the $L^{2}$-projection onto $\mathbb{P}_{k}(\mathcal{T}_{h})$, or its appropriate vector- or tensor-valued version.
+We denote the element-wise gradient and divergence operators by $\nabla_{h}$ (or ${\rm grad\,}_{h}$ for vectors) and $\operatorname{div}_{h}$, respectively.
+The main discrete spaces, operators, and interpolations are summarized in Table 1 for ease of reference.
+Table 1. Summary of main discrete spaces, operators, and interpolation maps.
+Object Description Role / Mapping
+$\mathring{\mathbb{V}}^{\operatorname{div}}_{k,\ell}(\mathcal{T}_{h})$ $H(\operatorname{div})$-conforming space Discrete velocity
+$\mathbb{P}_{\ell}^{-1}(\mathcal{T}_{h})$ Discontinuous polynomials Discrete pressure
+$\Sigma^{\rm tn}_{k}(\mathcal{T}_{h})$ t–n continuous traceless tensors Discrete pseudo-stress
+$\mathring{M}^{-1}_{k-1,k}(\mathbb{R}^{d})$ Vector-valued broken pair $\mathbb{P}_{k-1}(\mathcal{T}_{h};\mathbb{R}^{d})\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d})$
+$\operatorname{div}_{w}$ Weak divergence $\operatorname{div}_{w}:\Sigma^{\rm tn}_{k}\to\mathring{\mathbb{V}}^{\operatorname{div}}_{k,\ell}$
+$\operatorname{dev}{\rm grad\,}_{w}$ Weak deviatoric gradient $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}^{\operatorname{div}}_{k,\ell}\to\Sigma^{\rm tn}_{k}$
+$I^{\rm tn}_{k}$ t–n interpolant $I^{\rm tn}_{k}:H^{1}\to\Sigma^{\rm tn}_{k}$
+$I^{\operatorname{div}}_{k,\ell}$ $H(\operatorname{div})$ interpolant $I^{\operatorname{div}}_{k,\ell}:H^{1}\to\mathring{\mathbb{V}}^{\operatorname{div}}_{k,\ell}$
 2.2. Weak div stability for the scalar Laplacian
@@
 $\operatorname{div}:L^{2}(\Omega;\mathbb{R}^{d})\to H^{-1}(\Omega),$
-and discuss its distributional finite element discretization. This operator is part of the distributional de Rham complex introduced in [11].
+and study its distributional finite element discretization.
+This operator is part of the distributional de Rham complex introduced in [11].
+The scalar case allows the introduction of the mesh-dependent inner products and norms needed below, and it also reveals the main stability issues in distributional discretizations.
+This framework is then extended by a tensor-product construction in Section 2.3 to discretize the vector Laplacian, which is the main difficulty for $H(\operatorname{div})$-conforming velocity fields.
 The $L^{2}$ space is discretized by the discontinuous polynomial space
 $\Sigma_{k}^{-1}(\mathcal{T}_{h};\mathbb{R}^{d})=\mathbb{P}_{k}(\mathcal{T}_{h};\mathbb{R}^{d}),$
-which we abbreviate as $\Sigma_{k}^{-1}$. A natural norm on $\Sigma_{k}^{-1}$ is the standard $L^{2}$ norm $\|\cdot\|$.
+which we abbreviate as $\Sigma_{k}^{-1}$.
+The superscript 1-1 in $\Sigma_{k}^{-1}$ indicates that functions are taken elementwise with no interelement continuity imposed.
+A natural norm on $\Sigma_{k}^{-1}$ is the standard $L^{2}$ norm $\|\cdot\|$.
 Next, we introduce the product space of discontinuous polynomial spaces on $\mathcal{T}_{h}\times\mathcal{F}_{h}$:
-$M_{k-1,k}^{-1}:=\mathbb{P}_{k}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathcal{F}_{h}).$
-An element of $M_{k-1,k}^{-1}$ is denoted by $u=(u_{0},u_{b})$. A natural product-type inner product is defined by
+$M_{k-1,k}^{-1}:=\mathbb{P}_{k-1}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathcal{F}_{h}).$
+Again, the superscript 1-1 implies no continuity is imposed between the piecewise polynomial spaces.
+An element of $M_{k-1,k}^{-1}$ is denoted by $u=(u_{0},u_{b})$.
+A natural product-type inner product is defined by
 $((u_{0},u_{b}),(v_{0},v_{b}))_{0,h}=(u_{0},v_{0})+\sum_{F\in\mathcal{F}_{h}}h_{F}(u_{b},v_{b})_{F},$
@@
 The subspace
-$\mathring{M}_{k-1,k}^{-1}=\mathbb{P}_{k}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$
+$\mathring{M}_{k-1,k}^{-1}=\mathbb{P}_{k-1}(\mathcal{T}_{h})\times\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h})$
 can be viewed as a subspace of ${M}_{k-1,k}^{-1}$ by setting the values on boundary faces to zero.
 For $\sigma\in\Sigma_{k}^{-1}$, define the operator $\operatorname{div}_{w}:\Sigma_{k}^{-1}\to\mathring{M}_{k-1,k}^{-1}$ by
-$\operatorname{div}_{w}\sigma=\big(\operatorname{div}_{T}\sigma,-h_{F}^{-1}[\sigma\cdot\boldsymbol{n}]|_{F}\big)_{T\in\mathcal{T}_{h},~F\in\mathring{\mathcal{F}}_{h}}\in\mathring{M}_{k-1,k}^{-1}.$
-When $\sigma\in\Sigma_{k}^{-1}\cap H(\operatorname{div},\Omega)$, we have $\operatorname{div}_{w}\sigma=(\operatorname{div}\sigma,0)$ since $\sigma\cdot\boldsymbol{n}]|_{F}=0$ for all interior faces $F$; in this case we simply write $\operatorname{div}_{w}\sigma=\operatorname{div}\sigma$.
+$\operatorname{div}_{w}\sigma=\big(\operatorname{div}_{T}\sigma,-h_{F}^{-1}[\![\sigma\cdot\boldsymbol{n}]\!]|_{F}\big)_{T\in\mathcal{T}_{h},~F\in\mathring{\mathcal{F}}_{h}}\in\mathring{M}_{k-1,k}^{-1}.$
+When $\sigma\in\Sigma_{k}^{-1}\cap H(\operatorname{div},\Omega)$, we have $\operatorname{div}_{w}\sigma=(\operatorname{div}\sigma,0)$ since $\sigma\cdot\boldsymbol{n}]\!]|_{F}=0$ for all interior faces $F$; in this case we simply write $\operatorname{div}_{w}\sigma=\operatorname{div}\sigma$.
 A graph norm associated with $\operatorname{div}_{w}$ is defined as
-$\|\sigma\|_{\operatorname{div}_{w}}^{2}:=\|\sigma\|^{2}+\|\operatorname{div}_{w}\sigma\|_{0,h}^{2}=\|\sigma\|^{2}+\|\operatorname{div}_{h}\sigma\|^{2}+\|h^{-1/2}[\sigma\cdot\boldsymbol{n}]\|_{\mathring{\mathcal{F}}_{h}}^{2},$
+$\|\sigma\|_{\operatorname{div}_{w}}^{2}:=\|\sigma\|^{2}+\|\operatorname{div}_{w}\sigma\|_{0,h}^{2}=\|\sigma\|^{2}+\|\operatorname{div}_{h}\sigma\|^{2}+\|h^{-1/2}[\![\sigma\cdot\boldsymbol{n}]\!]\|_{\mathring{\mathcal{F}}_{h}}^{2},$
 where the scaling is chosen so that the last two terms are of comparable order.
@@
 Lemma 2.1.
-We have the weak div stability: there exists a constant $\alpha>0$, independent of $h$, such that
+We have the weak div stability: for a given $k\geq 0$, there exists a constant $\alpha>0$, independent of $h$, but might depend on $k$, such that
 (6) $\inf_{v\in\mathring{M}^{-1}_{k-1,k}}\sup_{\tau\in\Sigma_{k}^{-1}}\frac{(\operatorname{div}_{w}\tau,v)_{0,h}}{\|\tau\|_{\operatorname{div}_{w}}\|v\|_{0,h}}=\alpha>0.$
-Consider the mixed formulation of the Poisson equation $\Delta u=f$ with Dirichlet boundary condition $u|_{\partial\Omega}=0$.
-Find $\sigma\in\Sigma_{k}^{-1}$ and $u\in\mathring{M}_{k-1,k}^{-1}$ such that
-(4a) $(\sigma,\tau)$ $=(\nabla_{w}u,\tau)\quad\forall\,\tau\in\Sigma_{k}^{-1},$
-(4b) $(\operatorname{div}_{w}\sigma,v)$ $=-(f,v_{0})\quad\;\forall\,v\in\mathring{M}_{k-1,k}^{-1}.$
+Consider the mixed formulation of the Poisson equation $\Delta u=f$ with Dirichlet boundary condition $u|_{\partial\Omega}=0$: Find $\sigma\in\Sigma_{k}^{-1}$ and $u\in\mathring{M}_{k-1,k}^{-1}$ such that
+(7a) $(\sigma,\tau)+(\operatorname{div}_{w}\tau,u)_{0,h}$ $=0\qquad\qquad\;\forall\,\tau\in\Sigma_{k}^{-1},$
+(7b) $(\operatorname{div}_{w}\sigma,v)_{0,h}$ $=-(f,v_{0})\quad\;\forall\,v\in\mathring{M}_{k-1,k}^{-1}.$
 This formulation is precisely the hybridized mixed finite element method using the ${\rm BDM}_{k}$$\mathbb{P}_{k-1}$ pair [2]. The problem (7) is well-posed by the inf-sup condition (6).
@@
 In the Stokes problem, we use the weak divergence operator to discretize the vector Laplacian.
-Let $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ and $\Delta\boldsymbol{u}=-\operatorname{div}\boldsymbol{\sigma}$.
-To distinguish the vector formulation from the scalar case, we use boldface symbols such as $\boldsymbol{\sigma}$ and $\boldsymbol{u}$.
+Let $\boldsymbol{\sigma}=\nabla\boldsymbol{u}$; then, in the distributional sense, $\Delta\boldsymbol{u}=-\operatorname{div}\boldsymbol{\sigma}.$ To distinguish the vector formulation from the scalar case, we use boldface symbols such as $\boldsymbol{\sigma}$ and $\boldsymbol{u}$.
 By the tensor-product construction, we obtain the weak divergence operator
-$\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{M})\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}),$
+(8) $\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{M})\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}),$
 where
 $\Sigma_{k}^{-1}(\mathbb{X}):=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T;\mathbb{X})\quad\text{for }\mathbb{X}=\mathbb{M}\text{ or }\mathbb{T};\quad\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d}):=\mathbb{R}^{d}\otimes\mathring{M}_{k-1,k}^{-1}.$
-The tensor space is restricted to the traceless subspace $\mathbb{T}$, since for $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{div}\boldsymbol{u}=0$.
+The tensor space $\Sigma_{k}^{-1}(\mathbb{M})$ in (8) will be restricted to the traceless subspace $\Sigma_{k}^{-1}(\mathbb{T})$, since for $\boldsymbol{\sigma}={\rm grad\,}\boldsymbol{u}$ with $\operatorname{div}\boldsymbol{u}=0$, we have $\operatorname{tr}\boldsymbol{\sigma}=\operatorname{div}\boldsymbol{u}=0$.
 We then impose suitable continuity conditions on the range space and identify the subspaces of $\Sigma_{k}^{-1}(\mathbb{T})$ that lead to weakly divergence-stable pairs.
@@
 We employ $H(\operatorname{div})$-conforming finite elements to discretize the velocity field $\boldsymbol{u}$.
-For a simplex $T$ and integers $k\geq 0$ and $\ell\geq-1$, the local space of shape functions is defined by
-$\mathbb{V}_{k,\ell}^{\operatorname{div}}(T):=\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x},\quad\ell=k\text{ or }k-1.$
-For $\ell\geq 0$, the degrees of freedom (DoFs) are given by (cf. [16, Section 3])
+For a simplex $T$ and integers $k\geq 0$ and $\ell=k\text{ or }k-1$, the local space of shape functions is defined by
+$\mathbb{V}_{k,\ell}^{\operatorname{div}}(T):=\mathbb{P}_{k}(T;\mathbb{R}^{d})+\mathbb{H}_{\ell}(T)\boldsymbol{x}.$
+For $\ell\geq 0$, the degrees of freedom (DoFs) are given by (cf. [17, Section 3])
 (9a) $(\boldsymbol{v}\cdot\boldsymbol{n},q)_{F},$ $\quad q\in\mathbb{P}_{k}(F),\quad F\in\mathcal{F}(T),$
@@
 Lemma 2.2.
-The mapping $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ defined by
+For $k\geq 0$, and $\ell=k$ or k1k-1, the mapping $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ defined by
 $\boldsymbol{u}\mapsto\big(Q_{k-1,T}\boldsymbol{u},\,(\boldsymbol{0},\,\boldsymbol{n}_{F}\cdot\boldsymbol{u})_{F}\big)_{T\in\mathcal{T}_{h},\,F\in\mathring{\mathcal{F}}_{h}}$
@@
 Proof.
-If $(\boldsymbol{n}_{F}\cdot\boldsymbol{u})|_{F}=0$ and $Q_{k-1,T}\boldsymbol{u}=0$, then all degrees of freedom in (5) vanish, which implies $\boldsymbol{u}=0$. Hence, $E$ is injective. ∎
+If $(\boldsymbol{n}_{F}\cdot\boldsymbol{u})|_{F}=0$ and $Q_{k-1,T}\boldsymbol{u}=\boldsymbol{0}$, then all degrees of freedom in (9) vanish, which implies $\boldsymbol{u}=\boldsymbol{0}$.
+Hence, $E$ is injective.
+∎
+The tangential component $\Pi_{F}\boldsymbol{u}\in\mathbb{P}_{k}(F;\mathbb{R}^{d-1})$ is not necessarily continuous.
+We set this component to $\boldsymbol{0}$ here, as the stress space will satisfy tangential–normal continuity.
 The embedding $E$ can be extended to $H^{1}(\Omega;\mathbb{R}^{d})$, but the injectivity no longer holds.
 Spaces for the pseudo-stress
-For an integer $k\geq 0$, the local space of shape functions is $\mathbb{P}_{k}(T;\mathbb{T})$, with degrees of freedom (DoFs) defined by
+For an integer $k\geq 0$, the local space of shape functions is $\mathbb{P}_{k}(T;\mathbb{T})$, with DoFs defined by
 (10a) $\int_{F}\boldsymbol{t}_{i}^{\intercal}\boldsymbol{\tau}\boldsymbol{n}\,q\,\mathrm{d}S,$ $\quad q\in\mathbb{P}_{k}(F),\ i=1,2,\ldots,d-1,\ F\in\mathcal{F}(T),$
 (10b) $\int_{T}\boldsymbol{\tau}:\boldsymbol{q}\,\mathrm{d}x,$ $\quad\boldsymbol{q}\in\mathbb{P}_{k-1}(T;\mathbb{T}).$
-The unisolvence of these DoFs is established in [19, 30]. The corresponding global finite element space is
+The unisolvence of these DoFs for $\mathbb{P}_{k}(T;\mathbb{T})$ can be found in [20, 31]. The corresponding global finite element space is
 […]
-Functions in $\Sigma_{k}^{\rm tn}$ are tangential-normal continuous, i.e., $\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}]_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$.
+and we abbreviate $\Sigma_{k}^{\rm tn}(\mathcal{T}_{h})$ by $\Sigma_{k}^{\rm tn}$.
+Functions in $\Sigma_{k}^{\rm tn}$ are tangential-normal continuous, i.e., $\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}]\!]_{F}=0$ for all $F\in\mathring{\mathcal{F}}_{h}$.
 The weak divergence operator $\operatorname{div}_{w}$ defines a bilinear form on $\Sigma_{k}^{\rm tn}\times\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$:
-$(\operatorname{div}_{w}\boldsymbol{\sigma},\boldsymbol{v})_{0,h}=(\operatorname{div}_{w}\boldsymbol{\sigma},E\boldsymbol{v})_{0,h}$ $=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}],\boldsymbol{n}\cdot\boldsymbol{v})_{F}$
-$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}\boldsymbol{v})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}])_{F}.$
+$(\operatorname{div}_{w}\boldsymbol{\sigma},\boldsymbol{v})_{0,h}=(\operatorname{div}_{w}\boldsymbol{\sigma},E\boldsymbol{v})_{0,h}$ $=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\![\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}]\!],\boldsymbol{n}\cdot\boldsymbol{v})_{F}$
+$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\sigma},\operatorname{dev}{\rm grad\,}\boldsymbol{v})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n},[\![\Pi_{F}\boldsymbol{v}]\!])_{F}.$
 Remark 2.3.
-The bilinear form $(\operatorname{div}_{w}\cdot,\cdot)_{0,h}$ defines a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{\rm tn}\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})^{\prime}$.
-With the inner product $(\cdot,\cdot)_{0,h}$, we identify $\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ with its dual.
+The bilinear form $(\operatorname{div}_{w}\cdot,\cdot)_{0,h}$ induces a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{\rm tn}\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})^{\prime}$.
+With the inner product $(\cdot,\cdot)_{0,h}$, we can identify $\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ with its dual.
 Through the embedding $E$, we further identify $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cong E(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}})$ with its dual, and hence can write the mapping as
@@
 (11) $\|\boldsymbol{\tau}-I_{T}^{\rm tn}\boldsymbol{\tau}\|_{T}+h_{T}|\boldsymbol{\tau}-I_{T}^{\rm tn}\boldsymbol{\tau}|_{1,T}\lesssim h_{T}^{m}|\boldsymbol{\tau}|_{m,T}.$
-We now prove the following important commutative property.
+We now prove the following important commutative property which can be simply written as
+$Q_{k,k-1}^{\operatorname{div}}\operatorname{div}=\operatorname{div}_{w}I_{k}^{\rm tn},$
+where $Q_{k,k-1}^{\operatorname{div}}$ is the $L^{2}$-projection to the space $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$.
 Lemma 2.4.
@@
 Proof.
-Notice that $\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h}\in\Sigma_{k-1}^{-1}(\mathbb{T})$ and $\Pi_{F}\boldsymbol{v}_{h}]\in\mathbb{P}_{k}(F;\mathbb{R}^{d-1})$. So
-$(\operatorname{div}_{w}I_{k}^{\rm tn}\boldsymbol{\tau},\boldsymbol{v}_{h})_{0,h}$ $=-\sum_{T\in\mathcal{T}_{h}}(I_{k}^{\rm tn}\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}I_{k}^{\rm tn}\boldsymbol{\tau}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}_{h}])_{F}$
-$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\tau}\boldsymbol{n},[\Pi_{F}\boldsymbol{v}_{h}])_{F}$
+Notice that $\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}\in\Sigma_{k-1}^{-1}(\mathbb{T})$ and $\Pi_{F}\boldsymbol{v}_{h}]\!]|_{F}\in\mathbb{P}_{k}(F;\mathbb{R}^{d-1})$ for $F\in\mathcal{F}_{h}$.
+So
+$(\operatorname{div}_{w}I_{k}^{\rm tn}\boldsymbol{\tau},\boldsymbol{v}_{h})_{0,h}$ $=-\sum_{T\in\mathcal{T}_{h}}(I_{k}^{\rm tn}\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}(I_{k}^{\rm tn}\boldsymbol{\tau})\boldsymbol{n},[\![\Pi_{F}\boldsymbol{v}_{h}]\!])_{F}$
+$=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{\tau},\operatorname{dev}{\rm grad\,}\boldsymbol{v}_{h})_{T}+\sum_{F\in\mathcal{F}_{h}}(\Pi_{F}\boldsymbol{\tau}\boldsymbol{n},[\![\Pi_{F}\boldsymbol{v}_{h}]\!])_{F}$
 $=(\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}_{h}).$
@@
 Proof.
-For any $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\subset L^{2}(\Omega;\mathbb{R}^{d})$, there exists $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$ such that $\operatorname{div}\boldsymbol{\tau}=\boldsymbol{v}_{h}$.
-Let $\boldsymbol{\tau}_{h}=I_{k}^{\rm tn}\boldsymbol{\tau}$; then the commutative property (8) implies the desired result (9).
+For any $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\subset L^{2}(\Omega;\mathbb{R}^{d})$, the surjectivity of the divergence operator mapping from $H^{1}(\Omega;\mathbb{T})$ onto $L^{2}(\Omega;\mathbb{R}^{d})$ guarantees that there exists $\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T})$ such that $\operatorname{div}\boldsymbol{\tau}=\boldsymbol{v}_{h}$ and $\|\boldsymbol{\tau}\|_{1}\lesssim\|\boldsymbol{v}_{h}\|$ (cf.
+[6, (33)]).
+Let $\boldsymbol{\tau}_{h}=I_{k}^{\rm tn}\boldsymbol{\tau}$; then the commutative property (12) implies $\operatorname{div}_{w}\boldsymbol{\tau}_{h}=\boldsymbol{v}_{h}$, which, combined with the boundedness of the interpolant $I_{k}^{\rm tn}$, yields the desired result (13).
 ∎
-If the velocity space is enriched to ${\rm RT}_{k}$, i.e., $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$, then $\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})$.
+Remark 2.6.
+If the range space is enriched to ${\rm RT}_{k}$, i.e., $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$, then $\operatorname{dev}{\rm grad\,}_{h}\boldsymbol{v}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})$ not $\Sigma_{k-1}^{-1}(\mathbb{T})$.
 Since DoF (10b) to define $I_{k}^{\rm tn}$ is for moments of degree k1k-1, the commutative property (12) no longer holds.
-Consequently, the inf-sup condition may fail for the pair $\Sigma_{k}^{\rm tn}$-$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
+Consequently, the inf-sup condition (13) may fail for the pair $\Sigma_{k}^{\rm tn}$-$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
+Remark 2.7.
+On the other hand, to accommodate the slip boundary condition, we can reduce the stress space to impose the boundary condition $\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}=\boldsymbol{0}$ on $\partial\Omega$.
+Let
+$\mathring{H}^{1}(\Omega;\mathbb{T}):=\{\boldsymbol{\tau}\in H^{1}(\Omega;\mathbb{T}):\Pi_{F}\boldsymbol{\tau}\boldsymbol{n}=\boldsymbol{0}\text{ on }\partial\Omega\}.$
+Following the proof of (13), and using the surjectivity $\operatorname{div}\mathring{H}^{1}(\Omega;\mathbb{T})=L^{2}(\Omega;\mathbb{R}^{d})$ together with the commuting property (12), one can show that
+$\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}}\sup_{\boldsymbol{\tau}_{h}\in\mathring{\Sigma}_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|\boldsymbol{v}_{h}\|}=\alpha$
+for some constant $\alpha>0$ independent of $h$, where
+$\mathring{\Sigma}_{k}^{\rm tn}:=\{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}:\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}=\boldsymbol{0}\text{ on }\partial\Omega\}.$
 Define the weak deviatoric gradient $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\to\Sigma_{k}^{\rm tn}$ as the adjoint of $\operatorname{div}_{w}$.
@@
 Since $\operatorname{div}_{w}\Sigma_{k}^{\rm tn}=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, the operator $\operatorname{dev}{\rm grad\,}_{w}:\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\to\Sigma_{k}^{\rm tn}$ is injective, and thus $\|\operatorname{dev}{\rm grad\,}_{w}(\cdot)\|$ defines a norm on $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$.
-The inf-sup condition (9) can then be written equivalently as
+Consequently, the inf-sup condition (13) can be written equivalently as
 (14) $\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|\,\|\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{v}_{h}\|}=1>0.$
-Consider the mixed finite element method for the vector Laplacian
-$\Delta\boldsymbol{u}=\boldsymbol{f},\quad\text{ with }\;\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d}).$
-Find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ such that
-(11) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}$ $=0,\qquad\qquad\;\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},$
-$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.$
-The formulation (11) is well posed by the inf-sup conditions (9) or (10).
-However, the errors $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}I^{\operatorname{div}}_{k,k-1}\boldsymbol{u}\|$ are coupled.
-As a result, the dominant contribution is the $H^{1}$-type error of $\boldsymbol{u}$ of order $h^{k}$, rather than the $L^{2}$ error of $\boldsymbol{\sigma}$ of order $h^{k+1}$.
-This suboptimality comes from the loss of exact divergence-free conformity:
-$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0\ \ \forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\ \ \nRightarrow\ \ (\operatorname{div}_{w}\boldsymbol{\sigma}_{h},E\boldsymbol{v})_{0,h}=0\ \ \forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}).$
+With these preparations, we consider the mixed finite element method for the vector Laplacian: Find $\boldsymbol{u}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})$ such that $\Delta\boldsymbol{u}=\boldsymbol{f}$.
+The discrete problem seeks $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$ and $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ satisfying:
+(15) $\begin{cases}(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}=0&\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},\\ (\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=-(\boldsymbol{f},\boldsymbol{v}_{h})&\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.\end{cases}$
+The mixed method (15) is well-posed due to the inf-sup conditions (13) and (14).
+However, a standard error analysis reveals that $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}I^{\operatorname{div}}_{k,k-1}\boldsymbol{u}\|$ are coupled.
+Subtracting the continuous system from (15) yields:
+$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+(\boldsymbol{u}-\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}$ $=0,\quad\forall\,\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn},$
+$(\operatorname{div}_{w}(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}),\boldsymbol{v}_{h})_{0,h}$ $=0,\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}.$
+The term $(\boldsymbol{u}-\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}=(\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h},\boldsymbol{\tau}_{h})$ is bounded by the $H^{1}$-type error $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\|$ rather than by $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ alone.
+This suboptimality arises because being weakly divergence-free with respect to $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ does not imply strong divergence-free properties:
+$(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\quad\nRightarrow\quad\langle\operatorname{div}\boldsymbol{\sigma}_{h},\boldsymbol{v}\rangle=0\quad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}).$
+However, we will demonstrate in Lemma 2.9 that this property is preserved if the velocity test space in (15) is enriched from $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ to $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
+In this enriched setting, the test space is sufficiently large to ensure that for $\boldsymbol{\tau}_{h}=I_{h}^{\rm tn}\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}$, $\operatorname{div}_{w}\boldsymbol{\tau}_{h}=0$ holds.
+Consequently, the coupling term $(\boldsymbol{u}-\boldsymbol{u}_{h},\operatorname{div}_{w}\boldsymbol{\tau}_{h})_{0,h}=0$ vanishes, yielding the optimal stress estimate.
+By decoupling the stress error from the $O(h^{k})$ velocity approximation error, we avoid the lower-order contamination and achieve the full $O(h^{k+1})$ convergence rate.
 2.5. Commuting property for the weak dev grad operator
@@
 Proof.
-As $\operatorname{div}(\boldsymbol{\tau}_{h}|_{T})\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$ and $\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]|_{F}\in\mathbb{P}_{k}(F)$, we have
-$(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\tau}_{h})$ $=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(\boldsymbol{u}\cdot\boldsymbol{n},[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}])_{F},$
-$=-\sum_{T\in\mathcal{T}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}\cdot\boldsymbol{n},[\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}])_{F}$
+As $\operatorname{div}(\boldsymbol{\tau}_{h}|_{T})\in\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$ and $\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]\!]|_{F}\in\mathbb{P}_{k}(F)$, we have
+$(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\tau}_{h})$ $=-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(\boldsymbol{u}\cdot\boldsymbol{n},[\![\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]\!])_{F}$
+$=-\sum_{T\in\mathcal{T}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u},\operatorname{div}\boldsymbol{\tau}_{h})_{T}+\sum_{F\in\mathring{\mathcal{F}}_{h}}(I_{k,k}^{\operatorname{div}}\boldsymbol{u}\cdot\boldsymbol{n},[\![\boldsymbol{n}^{\intercal}\boldsymbol{\tau}_{h}\boldsymbol{n}]\!])_{F}$
 $=(\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}}\boldsymbol{u},\boldsymbol{\tau}_{h}).$
@@
 The weak operator $\operatorname{div}_{w}$ can be regarded as a discretization of $\operatorname{div}$ in the distributional sense.
-We can preserve the exact divergence condition in the following sense.
+We can preserve the strong divergence-free property in the following sense.
 Lemma 2.9.
@@
 $(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}=0,\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}},$
-then $\operatorname{div}\boldsymbol{\sigma}_{h}=0$.
+then $\operatorname{div}\boldsymbol{\sigma}_{h}=0$ in distribution sense.
 Proof.
@@
 2.6. Restriction to divergence-free subspaces
-The conditions (8) and (14) are incompatible.
-Weak div stability holds only on the smaller space $\mathbb{V}_{k,k-1}^{\operatorname{div}}$, while the exact divergence-freeness requires the larger space $\mathbb{V}_{k,k}^{\operatorname{div}}$.
-We resolve this conflict by considering the divergence-free subspaces.
-We have the following commuting property (cf. [5]):
-(15) $\operatorname{div}\!\big(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\big)=Q_{\ell}\big(\operatorname{div}\boldsymbol{v}\big),\qquad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}),$
-which implies $\operatorname{div}\!\big(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\big)=0$ for all $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$.
-The additional interior moments in ${\rm RT}_{k}$ compared with ${\rm BDM}_{k}$ serve only to enrich the range of the divergence operator.
-Hence, for any $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ and $k\geq 1$,
-$I_{k,k}^{\operatorname{div}}\boldsymbol{v}=I_{k,k-1}^{\operatorname{div}}\boldsymbol{v}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
+The commuting conditions (12) and (18) involve different velocity spaces, as shown in diagram (19):
+(19) $H^{1}(\Omega;\mathbb{T})$$L^{2}(\Omega;\mathbb{R}^{d})$$\Sigma_{k}^{\rm tn}$$\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$$\scriptstyle{\operatorname{div}}$$\scriptstyle{I_{k}^{\rm tn}}$$\scriptstyle{Q_{k,k-1}^{\operatorname{div}}}$$\scriptstyle{\operatorname{div}_{w}}$$H^{1}(\Omega;\mathbb{R}^{d})$$L^{2}(\Omega;\mathbb{T})$$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$$\Sigma_{k}^{\rm tn}$$\scriptstyle{\operatorname{dev}{\rm grad\,}}$$\scriptstyle{I_{k,k}^{\operatorname{div}}}$$\scriptstyle{Q_{k}^{\rm tn}}$$\scriptstyle{\operatorname{dev}{\rm grad\,}_{w}}$
+Weak divergence stability holds on the smaller space $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$, whereas strong divergence-freeness requires the larger space $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
+We resolve this structural mismatch by restricting our analysis to the divergence-free subspace.
+Recall the standard property for $H(\operatorname{div})$ interpolants [5]:
+(20) $\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v})=Q_{\ell}(\operatorname{div}\boldsymbol{v}),\quad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d}).$
+Equation (20) implies that $I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}\in\ker(\operatorname{div})$ whenever $\boldsymbol{v}\in\ker(\operatorname{div})$.
+Notably, for $k\geq 1$, the additional interior moments in $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$ compared to $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ only enrich the range of the divergence operator and do not alter the divergence-free part of the space.
+Consequently, for any $\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$, we have:
+(21) $I_{k,k}^{\operatorname{div}}\boldsymbol{v}=I_{k,k-1}^{\operatorname{div}}\boldsymbol{v}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
+This identity bridges the two diagrams, allowing us to leverage the stability of $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}$ alongside the consistency of the weak deviatoric gradient on $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$.
 Consider the mixed finite element method for the vector Laplacian
@@
 (23b) $(\operatorname{div}_{w}\boldsymbol{\sigma}_{h},\boldsymbol{v}_{h})_{0,h}$ $=-(\boldsymbol{f},\boldsymbol{v}_{h}),\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
+By the inf-sup condition (13) and the commuting property (18), we obtain the following two properties.
+Lemma 2.10.
+There exists a constant $\alpha>0$, independent of $h$, such that
+(24) $\inf_{\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})}\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{\rm tn}}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}\,\|\boldsymbol{v}_{h}\|}=\alpha.$
+Proof.
+It follows from the inf-sup condition (13) and the fact
+$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}\cap\ker(\operatorname{div})=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
+Lemma 2.11.
+We have the strong divergence-free property: if $(\operatorname{div}_{w}\boldsymbol{\tau}_{h},\boldsymbol{v}_{h})_{0,h}=0$ for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div})$, then in the distributional sense:
+$\langle\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v}\rangle=0\qquad\forall\,\boldsymbol{v}\in H^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div}).$
+Proof.
+By the definition of distributional $\operatorname{div}$, the commuting property (18), and (21)
+$\langle\operatorname{div}\boldsymbol{\tau}_{h},\boldsymbol{v}\rangle=-(\boldsymbol{\tau}_{h},\operatorname{dev}{\rm grad\,}\boldsymbol{v})=-(\boldsymbol{\tau}_{h},\operatorname{dev}{\rm grad\,}_{w}I_{k,k}^{\operatorname{div}}\boldsymbol{v})=-(\operatorname{div}_{w}\boldsymbol{\tau}_{h},I_{k,k}^{\operatorname{div}}\boldsymbol{v})=0.$
 Theorem 2.12.
-Let $k\geq 0$. The mixed finite element method (17) is well posed. Let $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u}$, where $\boldsymbol{u}$ solves (16). Then
+Let $k\geq 0$. The mixed finite element method (23) is well-posed. Let $\boldsymbol{\sigma}=\operatorname{dev}{\rm grad\,}\boldsymbol{u}$, where $\boldsymbol{u}$ solves (22). Then
 (25) $\|\operatorname{dev}{\rm grad\,}\boldsymbol{u}-\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\|=\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|\leq\|\boldsymbol{\sigma}-I_{k}^{\rm tn}\boldsymbol{\sigma}\|,$
@@
 Proof.
-The well-posedness follows from the inf-sup condition (9) and the fact
-$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}\cap\ker(\operatorname{div})=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
+The well-posedness is from (24).
 Denote by $\boldsymbol{\sigma}_{I}=I_{k}^{\rm tn}\boldsymbol{\sigma}$ and $\boldsymbol{u}_{I}=I_{k,k}^{\operatorname{div}}\boldsymbol{u}\in\mathbb{V}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div})$.
@@
 $(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{v}_{h})_{0,h}=0,\quad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\cap\ker(\operatorname{div}).$
-We verify the following orthogonality by using (14)
-$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})$ $=(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})+(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{h})_{0,h}$
-$=(\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{I},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})=-(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{I})_{0,h}=0.$
+So by Lemma 2.11, $(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})=0$. We obtain the following orthogonality
+$(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})$ $=(\operatorname{dev}{\rm grad\,}\boldsymbol{u},\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h})+(\operatorname{div}_{w}(\boldsymbol{\sigma}_{I}-\boldsymbol{\sigma}_{h}),\boldsymbol{u}_{h})_{0,h}=0.$
 The estimate (25) follows from this orthogonality:
@@
 ∎
-The velocity space $\boldsymbol{v}_{h}|_{T}\in\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$ and $\mathbb{P}_{k}(T;\mathbb{R}^{d})\subseteq\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, and the optimal order of convergence for $H^{1}$ type error is $h^{k}$.
-So (20) is a superconvergent error estimate for $\boldsymbol{u}$.
-3. Mixed Finite Element Methods for Stokes Equation
-In this section, we introduce the pressure as a Lagrange multiplier to enforce the divergence-free condition, and another multiplier to relax the tangential-normal continuity of the stress space.
-The resulting mixed finite element methods for the Stokes equation are pointwise divergence-free and superconvergent.
+Since $\mathbb{P}_{k}(T;\mathbb{R}^{d})\subseteq\mathbb{V}_{k,\ell}^{\operatorname{div}}(T)$, the standard approximation theory for the velocity space $\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ yields an $H^{1}$-type error of order $O(h^{k})$.
+Consequently, the estimate (27), which establishes a higher convergence rate $O(h^{k+1})$, constitutes a velocity superconvergence result.
+3. Mixed Finite Element Methods for the Stokes Equation
+In this section, we develop a mixed finite element discretization for the Stokes equation that ensures the pointwise divergence-free velocity.
+We introduce Lagrange multiplier spaces using piecewise polynomials for pressure and facial multipliers to relax the tangential-normal continuity of the pseudo-stress.
+After proving stability via a discrete inf-sup condition, we formulate the mixed scheme and show its equivalence to (23).
+The resulting error analysis yields optimal-order, pressure-robust bounds.
+Finally, we establish $L^{2}$-velocity superconvergence of order $O(h^{k+2})$ and present a postprocessing strategy that achieves $H^{1}$-superconvergence.
 3.1. Spaces of Lagrange multipliers
@@
 $\operatorname{div}\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}=\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}.$
-We introduce the Lagrange multiplier spaces
-$\Lambda_{k}=\mathbb{P}_{k}(\mathcal{F}_{h};\mathbb{R}^{d-1})\quad\textrm{and}\quad\mathring{\Lambda}_{k}=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1})$
-to relax the tangential-normal continuity condition in the space $\Sigma_{k}^{\rm tn}$.
-The embedding $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ is extended as
+To relax the tangential-normal continuity condition in the space $\Sigma_{k}^{\rm tn}$, we introduce the Lagrange multiplier spaces
+$\Lambda_{k}=\mathbb{P}_{k}(\mathcal{F}_{h};\mathbb{R}^{d-1})\quad\textrm{and}\quad\mathring{\Lambda}_{k}=\mathbb{P}_{k}(\mathring{\mathcal{F}}_{h};\mathbb{R}^{d-1}).$
+The embedding $E:\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}\to\mathring{M}_{k-1,k}^{-1}(\mathbb{R}^{d})$ is defined as
 $(\boldsymbol{u},\boldsymbol{\lambda})\mapsto\big(Q_{k-1,T}\boldsymbol{u},\,(\boldsymbol{\lambda},\,\boldsymbol{n}_{F}\cdot\boldsymbol{u})_{F}\big)_{T\in\mathcal{T}_{h},\,F\in\mathring{\mathcal{F}}_{h}},$
 which remains injective.
-The stress space is relaxed to $\Sigma_{k}^{-1}(\mathbb{T})=\prod_{T\in\mathcal{T}_{h}}\mathbb{P}_{k}(T;\mathbb{T})$.
+The stress space is relaxed to $\Sigma_{k}^{-1}(\mathbb{T})=\mathbb{P}_{k}(\mathcal{T}_{h};\mathbb{T})$.
 The weak divergence operator $\operatorname{div}_{w}$ then induces the bilinear form on $\Sigma_{k}^{-1}(\mathbb{T})\times(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$:
-(21) $(\operatorname{div}_{w}\boldsymbol{\sigma},(\boldsymbol{v},\boldsymbol{\mu}))_{0,h}=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\sum_{F\in\mathring{\mathcal{F}}_{h}}\!\Big(([\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}],\,\boldsymbol{n}\!\cdot\!\boldsymbol{v})_{F}+([\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}],\,\boldsymbol{\mu})_{F}\Big),$
+(28) $(\operatorname{div}_{w}\boldsymbol{\sigma},(\boldsymbol{v},\boldsymbol{\mu}))_{0,h}=\!\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v})_{T}-\!\sum_{F\in\mathring{\mathcal{F}}_{h}}\!\Big(([\![\boldsymbol{n}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}]\!],\,\boldsymbol{n}\cdot\boldsymbol{v})_{F}+([\![\Pi_{F}\boldsymbol{\sigma}\boldsymbol{n}]\!],\,\boldsymbol{\mu})_{F}\Big),$
 which induces a mapping, still denoted by $\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{T})\to(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})^{\prime}\cong(\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k})$ by the $(\cdot,\cdot)_{0,h}$ inner product through the embedding $E$.
@@
 Then (29) follows from the definition of the interpolation operator $I_{k,k}^{\operatorname{div}}$. ∎
+Again, (29) may not hold for $I_{k,k-1}^{\operatorname{div}}$, because the term $(I_{k,k}^{\operatorname{div}}\boldsymbol{v},\operatorname{div}\boldsymbol{\tau})_{T}$ requires the interior moments of the full space $\mathbb{P}_{k-1}(T;\mathbb{R}^{d})$.
+This property holds for RTk, but not for BDMk.
 3.2. Weak div stability
@@
 for each $T\in\mathcal{T}_{h}$, where $\boldsymbol{n}$ is the unit outward normal on $\partial T$.
-Then $\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}]=h_{F}\boldsymbol{\mu}_{h}$.
+Then $\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}]\!]=h_{F}\boldsymbol{\mu}_{h}$.
 By inverse and scaling estimates,
@@
 By the definition of $\operatorname{div}_{w}$,
-$(\operatorname{div}_{w}\boldsymbol{\tau}_{1},(0,\boldsymbol{\mu}_{h}))_{0,h}=-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}],\boldsymbol{\mu}_{h})_{F}=\|(0,\boldsymbol{\mu}_{h})\|_{0,h}^{2},$
+$(\operatorname{div}_{w}\boldsymbol{\tau}_{1},(0,\boldsymbol{\mu}_{h}))_{0,h}=-\sum_{F\in\mathring{\mathcal{F}}_{h}}([\![\Pi_{F}\boldsymbol{\tau}_{1}\boldsymbol{n}]\!],\boldsymbol{\mu}_{h})_{F}=\|(0,\boldsymbol{\mu}_{h})\|_{0,h}^{2},$
 and hence
 $\|(0,\boldsymbol{\mu}_{h})\|_{0,h}\lesssim\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},(0,\boldsymbol{\mu}_{h}))_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}}\qquad\forall\,\boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}.$
-On the other hand, by the weak divergence inf-sup (cf. (9)),
+On the other hand, by the weak divergence inf-sup condition (cf. (13)),
 $\|(\boldsymbol{v}_{h},0)\|_{0,h}=\|\boldsymbol{v}_{h}\|\lesssim\sup_{\boldsymbol{\tau}_{h}\in\Sigma_{k}^{-1}(\mathbb{T})}\frac{(\operatorname{div}_{w}\boldsymbol{\tau}_{h},(\boldsymbol{v}_{h},\boldsymbol{\mu}_{h}))_{0,h}}{\|\boldsymbol{\tau}_{h}\|_{\operatorname{div}_{w}}}\qquad\forall\,\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}},\ \boldsymbol{\mu}_{h}\in\mathring{\Lambda}_{k}.$
@@
 $\operatorname{div}_{w}:\Sigma_{k}^{-1}(\mathbb{T})\to\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}\times\mathring{\Lambda}_{k}.$
-The inf-sup condition (23) implies $\operatorname{div}_{w}$ is surjective. Consequently the adjoint $\operatorname{dev}{\rm grad\,}_{w}$ is injective and
-$\|\operatorname{dev}{\rm grad\,}_{w}(\cdot,\cdot)\|\text{ defines a norm on }\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\times\mathring{\Lambda}_{k}.$
-Notice that the velocity space cannot be enlarged to $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$. But later on we will use the coercivity on the divergence free subspaces.
+The inf–sup condition (30) ensures that the weak divergence operator $\operatorname{div}_{w}$ is surjective when $\ell=k-1$.
+Consequently, its adjoint, the weak deviatoric gradient $\operatorname{dev}{\rm grad\,}_{w}$, is injective, and the mapping
+$\|\operatorname{dev}{\rm grad\,}_{w}(\cdot,\cdot)\|$
+defines a norm on the product space $\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\times\mathring{\Lambda}_{k}$.
+It is important to note that the inf-sup condition (30) may not hold for the larger space $\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}$, even though this space is important for its approximation properties.
+In the subsequent analysis, the coercivity of the weak operators on the divergence-free subspaces is used to resolve this conflict.
 3.3. Mixed finite element methods
@@
 $\|Q_{\ell}p-p_{h}\|\lesssim\sup_{\boldsymbol{v}\in H_{0}^{1}(\Omega;\mathbb{R}^{d})}\frac{(\operatorname{div}(I_{k,\ell}^{\operatorname{div}}\boldsymbol{v}),p-p_{h})}{|\boldsymbol{v}|_{1}}\lesssim h^{k+1}|\boldsymbol{u}|_{k+2}.$
-The estimate (27) demonstrates that the mixed finite element method (24) is pressure-robust.
+The estimate (34) demonstrates that the mixed finite element method (31) is pressure-robust as it depends only on the regularity of $\boldsymbol{u}$ not pp.
 3.5. $L^{2}$ error estimate
@@
 (36) $\|\tilde{\boldsymbol{u}}\|_{2}\lesssim\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|.$
-We refer to [29, Remark I.5.6] and [44, Section 11.5] for the $H^{2}$-regularity (29) in convex domains.
+We refer to [30, Remark I.5.6] and [46, Section 11.5] for the $H^{2}$-regularity (36) in convex domains.
 Lemma 3.5.
@@
 $(\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h},{\rm grad\,}\tilde{\boldsymbol{u}})\lesssim h\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}|\tilde{\boldsymbol{u}}|_{2}.$
-Next consider case $k=0$. Applying the integration by parts to the right hand side of (31) to have
+Next consider case $k=0$. Applying integration by parts to the right hand side of (38) to have
 […]
 Since $\operatorname{div}\tilde{\boldsymbol{u}}=0$ and $\operatorname{div}H_{0}^{2}(\Omega;\mathbb{K})=H_{0}^{1}(\Omega;\mathbb{R}^{d})\cap\ker(\operatorname{div})$ (cf.
-[24, Theorem 1.1]), there exists a $\tilde{\boldsymbol{\tau}}\in H_{0}^{2}(\Omega;\mathbb{K})$ such that
+[25, Theorem 1.1]), there exists a $\tilde{\boldsymbol{\tau}}\in H_{0}^{2}(\Omega;\mathbb{K})$ such that
 $\operatorname{div}\tilde{\boldsymbol{\tau}}=\tilde{\boldsymbol{u}},\quad\textrm{and}\quad\|\tilde{\boldsymbol{\tau}}\|_{2}\lesssim\|\tilde{\boldsymbol{u}}\|_{1}.$
@@
 $\mathbb{V}_{h}^{d-2}:=\{\boldsymbol{\tau}_{h}\in H(\operatorname{div},\Omega;\mathbb{K}):\boldsymbol{\tau}_{h}|_{T}\in\mathbb{P}_{1}(T;\mathbb{K})\;\;\textrm{ for }T\in\mathcal{T}_{h}\}.$
-The DoFs for space $\mathbb{V}_{h}^{d-2}$ are given in [13, (3.10)].
+The DoFs for space $\mathbb{V}_{h}^{d-2}$ are given in [14, (3.10)].
 Let $I_{h}^{d-2}:H^{2}(\Omega;\mathbb{K})\to\mathbb{V}_{h}^{d-2}$ be the nodal interpolation operator.
@@
 $\operatorname{div}(I_{h}^{d-2}\boldsymbol{\tau})=I_{0,0}^{\operatorname{div}}(\operatorname{div}\boldsymbol{\tau})\quad\forall~\boldsymbol{\tau}\in H^{2}(\Omega;\mathbb{K}).$
-Then, employing the integration by parts and (26),
+Then, employing integration by parts and (33),
 […]
@@
 We can construct a superconvergent approximation of $\boldsymbol{u}$ using the estimates (34) and (39).
-For each $T\in\mathcal{T}_{h}$, find $\boldsymbol{u}_{h}^{*}|_{T}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d})$ and $p_{h}^{*}\in\mathbb{P}_{k}(T)\cap L_{0}^{2}(T)$ such that
+For each $T\in\mathcal{T}_{h}$, find $\boldsymbol{u}_{h}^{*}|_{T}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d})$ and $p_{h}^{*}\in\mathbb{P}_{k}(T)/\mathbb{R}$ such that
 (40a) $(\boldsymbol{u}_{h}^{*}\cdot\boldsymbol{n},q)_{F}$ $=(\boldsymbol{u}_{h}\cdot\boldsymbol{n},q)_{F}\quad\;\;\;\forall\,q\in\mathbb{P}_{0}(F),\ F\in\mathcal{F}(T),$
 (40b) $({\rm grad\,}\boldsymbol{u}_{h}^{*},{\rm grad\,}\boldsymbol{v})_{T}+(\operatorname{div}\boldsymbol{v},p_{h}^{*})_{T}$ $=(\boldsymbol{\sigma}_{h},{\rm grad\,}\boldsymbol{v})_{T}\quad\forall\,\boldsymbol{v}\in\widetilde{\mathbb{P}}_{k+1}(T;\mathbb{R}^{d}),$
-(33c) $(\operatorname{div}\boldsymbol{u}_{h}^{*},q)_{T}$ $=0\qquad\qquad\qquad\;\forall\,q\in{\mathbb{P}}_{k}(T)\cap L_{0}^{2}(T),$
+(40c) $(\operatorname{div}\boldsymbol{u}_{h}^{*},q)_{T}$ $=0\qquad\qquad\qquad\;\forall\,q\in\mathbb{P}_{k}(T)/\mathbb{R},$
 where $\widetilde{\mathbb{P}}_{k+1}(T;\mathbb{R}^{d}):=\{\boldsymbol{v}_{h}\in\mathbb{P}_{k+1}(T;\mathbb{R}^{d}):Q_{0,F}(\boldsymbol{v}_{h}\cdot\boldsymbol{n})=0\ \ \forall\,F\subset\partial T\}$.
-The local problems (33) are well-posed. From (33a) we get
-$Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h}^{*})=Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h})=0;$
-combined with (33c), this yields $\operatorname{div}\boldsymbol{u}_{h}^{*}=0$ on each $T\in\mathcal{T}_{h}$.
-Thus $\boldsymbol{u}_{h}^{*}\in\mathbb{P}_{k+1}(\mathcal{T}_{h};\mathbb{R}^{d})$ is pointwise divergence-free elementwise.
+The local problems (40) are well-posed. From (40a), it follows that
+$Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h}^{*})=Q_{0,T}(\operatorname{div}\boldsymbol{u}_{h})=0.$
+Combined with (40c), this identity implies that $\operatorname{div}\boldsymbol{u}_{h}^{*}=0$ for each $T\in\mathcal{T}_{h}$.
+Consequently, the postprocessed velocity $\boldsymbol{u}_{h}^{*}\in\mathbb{P}_{k+1}(\mathcal{T}_{h};\mathbb{R}^{d})$ is pointwise divergence-free on every element.
+However, we note that $\boldsymbol{u}_{h}^{*}$ is not globally $H(\operatorname{div})$-conforming, as the normal trace continuity (40a) is weakly enforced for $q\in\mathbb{P}_{0}(F)$ only.
 Theorem 3.7.
@@
 Combining (16), (39), and (34) gives (42). ∎
+Remark 3.8.
+The estimate for $\|Q_{\ell}p-p_{h}\|$ in (34) is optimal when $\ell=k$ and superconvergent when $\ell=k-1$. For fixed $k\geq 0$, since
+$\mathring{\mathbb{V}}_{k,k}^{\operatorname{div}}\cap\ker(\operatorname{div})=\mathring{\mathbb{V}}_{k,k-1}^{\operatorname{div}}\cap\ker(\operatorname{div}),$
+the discrete solutions $\boldsymbol{u}_{h}$ and $\boldsymbol{\lambda}_{h}$ are independent of $\ell$.
+In fact, if $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},p_{h})$ solves the mixed method (31) with $\ell=k$, then $(\boldsymbol{u}_{h},\boldsymbol{\lambda}_{h},Q_{k-1}p_{h})$ is the solution of the method with $\ell=k-1$.
+When $k\geq 1$, we have $p_{h}-Q_{k-1}p_{h}\in\mathbb{P}_{k}(T)/\mathbb{R}$ and
+$\operatorname{div}\big(\mathbb{V}_{k,k}^{\operatorname{div}}(T)\cap H_{0}(\operatorname{div},T)\big)=\mathbb{P}_{k}(T)/\mathbb{R},$
+so $p_{h}$ can be recovered locally from $Q_{k-1}p_{h}$. Specifically, for each $T\in\mathcal{T}_{h}$, $p_{h}-Q_{k-1}p_{h}$ is characterized by
+$(\operatorname{div}\boldsymbol{v}_{h},p_{h}-Q_{k-1}p_{h})_{T}=(\boldsymbol{f},\boldsymbol{v}_{h})_{T}-(\boldsymbol{\sigma}_{h},\operatorname{dev}{\rm grad\,}_{w}(\boldsymbol{v}_{h},0))_{T}-(\operatorname{div}\boldsymbol{v}_{h},Q_{k-1}p_{h})_{T},$
+for all $\boldsymbol{v}_{h}\in\mathbb{V}_{k,k}^{\operatorname{div}}(T)\cap H_{0}(\operatorname{div},T)$.
+We refer to [9, 54, 41] for related pressure post-processing procedures in virtual element methods.
 4. Equivalent Discrete Methods
-In this section, we explore several discrete methods that are equivalent to the mixed finite element scheme (24).
+We establish connections between our mixed finite element method and several existing divergence-free schemes.
+We first show that hybridizing the velocity leads to an equivalent $H^{1}$-nonconforming virtual element formulation, facilitating implementation in low-order cases.
+Second, we demonstrate equivalence to stabilization-free virtual element methods, where the weak deviatoric gradient acts as a local projection operator.
+Finally, we compare our approach with the Mass-Conserving-Stress (MCS) method and show that low-order cases can be handled naturally within our framework.
 4.1. Hybridization of the velocity
-Recall the $H^{1}$-nonconforming virtual element in [7, 15]. For integers $k\geq 0$ and $\ell=k,k-1$, the shape function space is defined as
-$V_{k+1,\ell+2}^{\rm VE}(T)=\{v\in H^{1}(T):\Delta v\in\mathbb{P}_{\ell}(T),\partial_{n}v|_{F}\in\mathbb{P}_{k}(F)\textrm{ for }F\in\mathcal{F}(T)\}.$
+We recall the $H^{1}$-nonconforming virtual element in [8, 16]. For integers $k\geq 0$ and $\ell=k,k-1$, the shape function space is defined as
+(44) $V_{k+1,\ell+2}^{\rm VE}(T)=\{v\in H^{1}(T):\Delta v\in\mathbb{P}_{\ell}(T),\partial_{n}v|_{F}\in\mathbb{P}_{k}(F)\textrm{ for }F\in\mathcal{F}(T)\}.$
 The DoFs for $V_{k+1,\ell+2}^{\rm VE}(T)$ are given by
@@
 4.2. Stabilization-free virtual element methods
-For $k\geq 0$, define the space of shape functions for an $H^{1}$-nonconforming virtual element as
+For $k\geq 0$, define the space of vector-valued shape functions for an $H^{1}$-nonconforming virtual element as
 $\mathbb{V}_{k+1}^{\rm VE}(T):=\{\boldsymbol{v}\in H^{1}(T;\mathbb{R}^{d}):$ $\operatorname{div}\boldsymbol{v}\in\mathbb{P}_{k}(T),\textrm{ there exists some }s\in L^{2}(T)$
@@
 $\textrm{ and }(\partial_{n}\boldsymbol{v}+s\boldsymbol{n})|_{F}\in\mathbb{P}_{k}(F;\mathbb{R}^{d})\;\forall~F\in\mathcal{F}(T)\}.$
-Following the argument in [53, Section 3.1], the local virtual element space $\mathbb{V}_{k+1}^{\rm VE}(T)$ is uniquely determined by the following DoFs
+Following the argument in [54, Section 3.1], the local virtual element space $\mathbb{V}_{k+1}^{\rm VE}(T)$ is uniquely determined by the following DoFs
 (48a) $(\boldsymbol{v},\boldsymbol{q})_{F},$ $\quad\boldsymbol{q}\in\mathbb{P}_{k}(F;\mathbb{R}^{d}),F\in\mathcal{F}(T),$
@@
 for all $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{\rm VE}$ and $q_{h}\in\mathbb{P}_{k}(\mathcal{T}_{h})/\mathbb{R}$.
-When $k=0$, the mixed method (41) is exactly the modified Crouzeix-Raviart element method in [42, (23)].
+When $k=0$, the mixed method (49) is exactly the modified Crouzeix-Raviart element method in [44, (23)].
 Lemma 4.2.
@@
 Remark 4.4.
-We can also use the following virtual element space for the velocity:
+We can also use the following virtual element space for the velocity based on (44):
 […]
 […]
-4.3. Discrete methods based on the pseudostress-velocity-pressure formulation
+4.3. Mass-Conserving-Stress Method
 By introducing $\boldsymbol{\sigma}_{h}:=\operatorname{dev}{\rm grad\,}_{w}\boldsymbol{u}_{h}\in\Sigma_{k}^{\rm tn}$, we can rewrite the mixed finite element method (31) as the MCS method [32]: find $\boldsymbol{\sigma}_{h}\in\Sigma_{k}^{\rm tn}$, $\boldsymbol{u}_{h}\in\mathring{\mathbb{V}}_{k,\ell}^{\operatorname{div}}$ and $p_{h}\in\mathbb{P}_{\ell}(\mathcal{T}_{h})/\mathbb{R}$ such that
@@
 was employed in the MCS method [32] to discretize the traceless stress tensor $\boldsymbol{\sigma}$.
-However, the constraint $\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{d-1})$ prevents the discrete scheme in [31] from achieving superconvergence properties such as the estimates $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|Q_{\ell}p-p_{h}\|$ in (27), and $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|$ in (32).
+However, the constraint $\Pi_{F}\boldsymbol{\tau}_{h}\boldsymbol{n}\in\mathbb{P}_{k-1}(F;\mathbb{R}^{d-1})$ prevents the decoupling of the error and thus not achieving superconvergence properties such as the estimates $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ and $\|Q_{\ell}p-p_{h}\|$ in (34), and $\|I_{k,k}^{\operatorname{div}}\boldsymbol{u}-\boldsymbol{u}_{h}\|$ in (39).
 Superconvergent results were later obtained in [31] by enriching the stress space, but their approach requires $k=\ell\geq 1$, whereas our formulation also covers low-order cases: $(k,\ell)=(0,-1)$, $(0,0)$, and $(1,0)$.
@@
 5. Numerical Examples
-In this section, we present numerical examples to validate the error estimates of the mixed finite element method (24) and the post-processing procedure (33).
-The test problems are adapted from [30], which primarily considered higher-order cases with $k=\ell\geq 1$.
-Here, we focus on low-order settings with $\ell=0$ and $k=0,1$.
-The computational domain is $\Omega=(0,1)^{d}$ for $d=2,3$, uniformly partitioned into simplicial meshes. All numerical experiments were developed based on the MATLAB package iiFEM [14].
+In this section, we verify the error estimates for the mixed finite element method (31) and the postprocessing scheme (40).
+Both smooth and singular test problems are considered, with focus on the low-order cases $\ell=0$ and $k=0,1$.
+The smooth benchmark problems are adapted from [31], where examples for $k=\ell\geq 1$ are presented.
+Throughout, the computational domains are discretized by uniform simplicial meshes, and all computations are carried out in MATLAB using iiFEM [15].
+5.1. Smooth problems with non-slip boundary conditions
 Example 5.1.
-We first test the mixed finite element method (24) in two dimensions. Take the exact solution
-$\boldsymbol{u}=\operatorname{curl}\psi_{2},\quad\text{and}\quad p=-x^{5}-y^{5}+\frac{1}{3},$
+We first test the mixed finite element method (31) in two dimensions on the unit square domain $\Omega=(0,1)^{2}$. The exact solution is
+$\boldsymbol{u}=\operatorname{curl}\psi_{2},\qquad p=-x^{5}-y^{5}+\frac{1}{3},$
 where $\psi_{2}=x^{2}(x-1)^{2}y^{2}(y-1)^{2}$.
-Table 1. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, and $\|p-p_{h}\|$ for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
-$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ order $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ order $\|p-p_{h}\|$ order
-$2^{-3}$ $(0,0)$ 2.988e-03 3.103e-02 7.810e-02
-$2^{-4}$ $(0,0)$ 1.284e-03 1.22 1.677e-02 0.89 3.914e-02 1.00
-$2^{-5}$ $(0,0)$ 5.988e-04 1.10 8.700e-03 0.95 1.963e-02 1.00
-$2^{-6}$ $(0,0)$ 2.907e-04 1.04 4.440e-03 0.97 9.840e-03 1.00
-$2^{-7}$ $(0,0)$ 1.436e-04 1.02 2.247e-03 0.98 4.931e-03 1.00
-$2^{-3}$ $(1,0)$ 3.296e-04 2.447e-03 7.453e-02
-$2^{-4}$ $(1,0)$ 8.382e-05 1.98 6.305e-04 1.96 3.760e-02 0.99
-$2^{-5}$ $(1,0)$ 2.104e-05 1.99 1.597e-04 1.98 1.880e-02 1.00
-$2^{-6}$ $(1,0)$ 5.266e-06 2.00 4.016e-05 1.99 9.428e-03 1.00
-$2^{-7}$ $(1,0)$ 1.317e-06 2.00 1.007e-05 2.00 4.715e-03 1.00
-Table 2. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
-$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
-$2^{-3}$ $(0,0)$ 1.233e-03 2.890e-02
-$2^{-4}$ $(0,0)$ 3.277e-04 1.91 1.481e-02 0.96
-$2^{-5}$ $(0,0)$ 8.353e-05 1.97 7.453e-03 0.99
-$2^{-6}$ $(0,0)$ 2.099e-05 1.99 3.733e-03 1.00
-$2^{-7}$ $(0,0)$ 5.256e-06 2.00 1.867e-03 1.00
-$2^{-3}$ $(1,0)$ 3.296e-05 2.286e-03
-$2^{-4}$ $(1,0)$ 4.167e-06 2.98 5.183e-04 1.98
-$2^{-5}$ $(1,0)$ 5.264e-07 2.99 1.463e-04 1.99
-$2^{-6}$ $(1,0)$ 6.625e-08 2.99 3.666e-05 2.00
-$2^{-7}$ $(1,0)$ 8.315e-09 2.99 9.178e-06 2.00
+Table 2. Errors for Example 5.1 with $\ell=0$ and $k=0,1$ in two dimensions.
+$h$ $(k,\ell)$ $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ order $\|p-p_{h}\|$ order $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|\nabla_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
+$2^{-3}$ $(0,0)$ 3.103e-02 7.810e-02 1.233e-03 2.890e-02
+$2^{-4}$ $(0,0)$ 1.677e-02 0.89 3.914e-02 1.00 3.277e-04 1.91 1.481e-02 0.96
+$2^{-5}$ $(0,0)$ 8.700e-03 0.95 1.963e-02 1.00 8.353e-05 1.97 7.453e-03 0.99
+$2^{-6}$ $(0,0)$ 4.440e-03 0.97 9.840e-03 1.00 2.099e-05 1.99 3.733e-03 1.00
+$2^{-7}$ $(0,0)$ 2.247e-03 0.98 4.931e-03 1.00 5.256e-06 2.00 1.867e-03 1.00
+$2^{-3}$ $(1,0)$ 2.447e-03 7.453e-02 3.296e-05 2.286e-03
+$2^{-4}$ $(1,0)$ 6.305e-04 1.96 3.760e-02 0.99 4.167e-06 2.98 5.183e-04 1.98
+$2^{-5}$ $(1,0)$ 1.597e-04 1.98 1.880e-02 1.00 5.264e-07 2.99 1.463e-04 1.99
+$2^{-6}$ $(1,0)$ 4.016e-05 1.99 9.428e-03 1.00 6.625e-08 2.99 3.666e-05 2.00
+$2^{-7}$ $(1,0)$ 1.007e-05 2.00 4.715e-03 1.00 8.315e-09 2.99 9.178e-06 2.00
 Example 5.2.
-We then test the mixed finite element method (24) in three dimensions. Take the exact solution
-$\boldsymbol{u}=\operatorname{curl}(\psi_{3},\psi_{3},\psi_{3})^{\intercal},\quad\text{and}\quad p=-x^{5}-y^{5}-z^{5}+\frac{1}{2},$
+We next consider the mixed finite element method (31) in three dimensions on the unit cube domain $\Omega=(0,1)^{3}$. The exact solution is given by
+$\boldsymbol{u}=\operatorname{curl}(\psi_{3},\psi_{3},\psi_{3})^{\intercal},\qquad p=-x^{5}-y^{5}-z^{5}+\frac{1}{2},$
 where $\psi_{3}=x^{2}(x-1)^{2}y^{2}(y-1)^{2}z^{2}(z-1)^{2}$.
-The numerical errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, and $\|p-p_{h}\|$ are reported in Table 1 for Example 5.1 and Table 3 for Example 5.2, while the post-processed errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ are given in Table 2 for Example 5.1 and Table 4 for Example 5.2.
-The results in these tables demonstrate that
-$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|\eqsim\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|=\mathcal{O}(h^{k+1}),\quad\|p-p_{h}\|=\mathcal{O}(h),$
-and for the post-processed solution $\boldsymbol{u}_{h}^{*}$,
-$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|=\mathcal{O}(h^{k+2}),\quad\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|=\mathcal{O}(h^{k+1}),$
-which are consistent with the theoretical estimates (27) and (34)–(35).
-Table 3. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$, $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$, and $\|p-p_{h}\|$ for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
+Table 3. Errors for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
+$h$ $(k,\ell)$ $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ order $\|p-p_{h}\|$ order $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|\nabla_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
+$2^{-1}$ $(0,0)$ 9.787e-03 2.942e-01 4.167e-04 4.576e-03
+$2^{-2}$ $(0,0)$ 4.960e-03 1.06 1.649e-01 0.84 1.565e-04 1.41 2.880e-03 0.67
+$2^{-3}$ $(0,0)$ 2.431e-03 0.95 8.501e-02 0.96 4.400e-05 1.83 1.539e-03 0.90
+$2^{-4}$ $(0,0)$ 1.260e-03 0.95 4.284e-02 0.99 1.143e-05 1.95 7.836e-04 0.97
+$2^{-5}$ $(0,0)$ 6.397e-04 0.98 2.146e-02 1.00 2.889e-06 1.98 3.937e-04 0.99
+$2^{-1}$ $(1,0)$ 2.431e-03 2.942e-01 9.624e-05 1.894e-03
+$2^{-2}$ $(1,0)$ 5.715e-04 2.09 1.649e-01 0.84 1.493e-05 2.69 5.478e-04 1.79
+$2^{-3}$ $(1,0)$ 1.541e-04 1.89 8.501e-02 0.96 2.035e-06 2.88 1.486e-04 1.88
+$2^{-4}$ $(1,0)$ 3.975e-05 1.96 4.284e-02 0.99 2.627e-07 2.95 3.808e-05 1.96
+Numerical errors for the smooth problems (Examples 5.1 and 5.2) are reported in Tables 2 and 3. The observed rates agree with the theoretical estimates (34) and (41)–(42). In particular,
+$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}=\mathcal{O}(h^{k+1}),\qquad\|p-p_{h}\|=\mathcal{O}(h),$
+and the postprocessed velocity satisfies
+$\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|=\mathcal{O}(h^{k+2}),\qquad\|\nabla_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|=\mathcal{O}(h^{k+1}).$
+5.2. Singular problem with slip boundary conditions
+We next consider a singular test problem on the two-dimensional L-shaped domain $\Omega=(-1,1)^{2}\setminus\big([0,1]\times[-1,0]\big)$ with boundary condition
+$\boldsymbol{u}\cdot\boldsymbol{n}=0\quad\text{and}\quad\boldsymbol{t}^{\intercal}\boldsymbol{\sigma}\boldsymbol{n}=0\quad\text{on }\partial\Omega.$
+The re-entrant corner at the origin induces reduced regularity of the exact solution and hence reduced convergence rates on uniform meshes.
+We impose this boundary condition by using the space $\mathring{\Sigma}_{k}^{\rm tn}$. See Remark 2.7.
+Example 5.3.
+Let $(r,\theta)$ denote polar coordinates centered at the re-entrant corner of $\Omega$. The corner is located at $(0,0)$, and $\theta\in(0,3\pi/2)$ in $\Omega$. Define,
+$\Phi(x,y)=(1-x^{2})^{3}(1-y^{2})^{3},\qquad\Psi_{s}(r,\theta)=r^{1+\lambda}\sin((\lambda-1)\theta),\qquad\lambda=\tfrac{1}{3},$
+and set
+$\boldsymbol{u}={\rm curl\,}(\Phi\Psi_{s}).$
+Further, let
+$g(\theta)=\sin((\lambda-1)\theta),\quad Q(\theta)=(1+\lambda)^{2}g(\theta)+g^{\prime\prime}(\theta),\quad p_{s}(r,\theta)=\frac{1}{\lambda-1}r^{\lambda-1}Q^{\prime}(\theta).$
+The pressure is defined by
+$p(x,y)=\Phi(x,y)p_{s}(r,\theta)-c_{0},\quad c_{0}=|\Omega|^{-1}\int_{\Omega}\Phi p_{s}\,dx,$
+so that $p\in L_{0}^{2}(\Omega)$.
+(a) $\boldsymbol{u}_{h}\text{ for }k=0,\ell=0$
+(b) $\boldsymbol{u}_{h}\text{ for }k=1,\ell=0$
+(c) Exact solution $\boldsymbol{u}$
+Figure 2. Flow structure of the velocity field $\boldsymbol{u}$ on the L-shaped domain. Left and middle: numerical solutions. Right: exact solution.
+Table 4. Errors for Example 5.3 with $\ell=0$ and $k=0,1$ in the L-shaped domain.
 $h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ order $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}$ order $\|p-p_{h}\|$ order
-$2^{-1}$ $(0,0)$ 6.762e-04 9.787e-03 2.942e-01
-$2^{-2}$ $(0,0)$ 3.110e-04 1.12 4.960e-03 1.06 1.649e-01 0.84
-$2^{-3}$ $(0,0)$ 1.399e-04 1.15 2.431e-03 0.95 8.501e-02 0.96
-$2^{-4}$ $(0,0)$ 6.719e-05 1.06 1.260e-03 0.95 4.284e-02 0.99
-$2^{-5}$ $(0,0)$ 3.320e-05 1.02 6.397e-04 0.98 2.146e-02 1.00
-$2^{-1}$ $(1,0)$ 2.076e-04 2.431e-03 2.942e-01
-$2^{-2}$ $(1,0)$ 6.700e-05 1.63 5.715e-04 2.09 1.649e-01 0.84
-$2^{-3}$ $(1,0)$ 1.790e-05 1.90 1.541e-04 1.89 8.501e-02 0.96
-$2^{-4}$ $(1,0)$ 4.551e-06 1.98 3.975e-05 1.96 4.284e-02 0.99
-Table 4. Errors $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ and $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ for Example 5.2 with $\ell=0$ and $k=0,1$ in three dimensions.
-$h$ $(k,\ell)$ $\|\boldsymbol{u}-\boldsymbol{u}_{h}^{*}\|$ order $\|{\rm grad\,}_{h}(\boldsymbol{u}-\boldsymbol{u}_{h}^{*})\|$ order
-$2^{-1}$ $(0,0)$ 4.167e-04 4.576e-03
-$2^{-2}$ $(0,0)$ 1.565e-04 1.41 2.880e-03 0.67
-$2^{-3}$ $(0,0)$ 4.400e-05 1.83 1.539e-03 0.90
-$2^{-4}$ $(0,0)$ 1.143e-05 1.95 7.836e-04 0.97
-$2^{-5}$ $(0,0)$ 2.889e-06 1.98 3.937e-04 0.99
-$2^{-1}$ $(1,0)$ 9.624e-05 1.894e-03
-$2^{-2}$ $(1,0)$ 1.493e-05 2.69 5.478e-04 1.79
-$2^{-3}$ $(1,0)$ 2.035e-06 2.88 1.486e-04 1.88
-$2^{-4}$ $(1,0)$ 2.627e-07 2.95 3.808e-05 1.96
+$1/14$ $(0,0)$ 1.866e-01 1.456e+00 1.364e+00
+$1/28$ $(0,0)$ 1.156e-01 0.70 1.023e+00 0.51 1.006e+00 0.44
+$1/56$ $(0,0)$ 7.488e-02 0.63 7.737e-01 0.40 7.250e-01 0.47
+$1/112$ $(0,0)$ 4.874e-02 0.62 6.093e-01 0.35 5.170e-01 0.49
+$1/224$ $(0,0)$ 3.153e-02 0.63 4.865e-01 0.33 3.693e-01 0.49
+$1/14$ $(1,0)$ 7.867e-02 1.791e+00 8.350e-01
+$1/28$ $(1,0)$ 5.253e-02 0.59 1.396e+00 0.36 6.071e-01 0.46
+$1/56$ $(1,0)$ 3.455e-02 0.60 1.097e+00 0.35 4.423e-01 0.46
+$1/112$ $(1,0)$ 2.242e-02 0.63 8.638e-01 0.34 3.257e-01 0.44
+$1/224$ $(1,0)$ 1.440e-02 0.64 6.820e-01 0.34 2.434e-01 0.42
+The numerical results for Example 5.3 are reported in Table 4. The corner singularity implies that $\boldsymbol{\sigma},p\in H^{1/3-\varepsilon}(\Omega)$. Hence, on uniform meshes, one expects
+$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|_{0,h}=\mathcal{O}(h^{1/3}),\quad\|p-p_{h}\|=\mathcal{O}(h^{1/3}),\quad\|\boldsymbol{u}-\boldsymbol{u}_{h}\|=\mathcal{O}(h^{2/3}).$
+The observed rates are consistent with these predictions.
+Fig.
+2 shows the velocity field on the L-shaped domain for two numerical solutions and the exact solution.
+In particular, the effect of the re-entrant corner is clearly visible.
+Although the convergence order of $\boldsymbol{u}_{h}$ is limited by the regularity, BDM1 with $(k,\ell)=(1,0)$ still performs better than RT0 with $(k,\ell)=(0,0)$ away from the singularity.
+6. Conclusions and Future Work
+In this paper, a family of pointwise divergence-free mixed finite element methods is developed for the Stokes equation.
+By using a distributional discretization of the vector Laplacian, the method gives a consistent and stable scheme even though the velocity space is only $H(\operatorname{div})$-conforming.
+The central mechanism is the weak deviatoric gradient operator $\operatorname{dev}{\rm grad\,}_{w}$, which uses tangential–normal continuous traceless tensor elements to preserve discrete integration-by-parts identities and commuting properties.
+In addition, a supercloseness result is established for the velocity and pressure approximations, and a post-processed discrete velocity achieves super-convergent rate.
+The framework also points to several directions for future work:
+  1. (1)
+    Brinkman and Navier–Stokes equations: The $\operatorname{dev}{\rm grad\,}_{w}$ framework is well suited to the Brinkman model, where the weak deviatoric gradient operator can naturally handle the transition between Darcy and Stokes regimes.
+For high-Reynolds-number Navier–Stokes flows, additional stabilization is still needed when convection dominates.
+  2. (2)
+    Linear elasticity: The link between stress-like variables and divergence constraints suggests extensions to mixed elasticity.
+Symmetric tensor spaces in this framework may lead to robust elements for nearly incompressible materials.
+  3. (3)
+    Boundary conditions and regularity: The weak operators can be adapted to slip and traction boundary conditions through modified boundary trace terms.
+It is also important to study superconvergence and adaptive finite element methods on nonconvex domains.